### v1

# Cost-benefit analysis of an AI-driven operational digital platform for integrated electric mobility, renewable energy, and grid management

Abstract
Integrating electric mobility (electric vehicles (EVs), electric trucks (ETs)) and renewable energy sources (RES) with the power grid is paramount for achieving decarbonization, efficiency, and stability. Given the rapid growth of decentralized technologies and their critical role in decarbonization, two critical challenges emerge: first, the development of a digital platform for operational coordination; and second, rigorous research into their cost–benefit profile. This paper addresses this by presenting a comprehensive cost–benefit analysis (CBA) of an AI-driven operational digital platform (ODP) designed for holistic, cross-sectoral optimization. The ODP aims to enhance energy efficiency, grid reliability, and environmental sustainability. A seven-step CBA framework, aligned with EU guidelines, quantifies economic, reliability, and environmental benefits against capital and operational expenditures, explicitly linking benefit magnitude to AI-driven ODP and optimization efficiencies, such as quantified improvements in market arbitrage from ODP, enabled forecasting, and enhanced operational efficiencies across various services. The framework is illustrated with case studies from Austria, Hungary, and Slovenia. Results show a robust benefit-cost ratio (BCR) of approximately 1.41 and a net present value (NPV) exceeding €356 million over a ten-year operational period (2026–2035). Extensive sensitivity analyses, including discount-rate variability and Monte Carlo simulations on BCR and NPV, confirm the framework’s robustness. This research underscores AI-powered integration as critical for achieving decarbonization and efficiency goals in the evolving energy and transport sectors.
keywords:
Operational digital platform , Cost-benefit analysis , AI-driven optimization , Renewable energy integration , Transport electrification1 Introduction
The European Union (EU) is undergoing a profound transformation towards sustainable energy and transport systems, characterized by the accelerated adoption of electric vehicles (EVs), electric trucks (ETs), and an increasing share of renewable energy sources (RES) in the power grid [1, 2, 3, 4]. The European Union has set the ambitious goal of achieving 60 % EV/ET sales of total vehicles by 2030 [5]. The effective integration of burgeoning EV/ET charging demands with intermittent RES generation and existing grid infrastructure presents substantial technical and economic challenges [6, 7]. These include managing the spatiotemporal mismatch between variable renewable supply and dynamic charging demand, enhancing grid flexibility to accommodate new, potentially high-impact loads, and overcoming the lack of coordinated platforms for holistic management across the energy and transport sectors. Traditional approaches often operate in silos, leading to suboptimal asset utilization, grid congestion, curtailment of renewable energy, and underutilization of emerging flexibility resources.
The challenge becomes more pronounced for heavy-duty vehicles. Studies have shown that battery-electric propulsion is technically viable but economically constrained for long-haul freight applications due to battery weight, range limitations, and charging infrastructure gaps [8, 9]. Additionally, analyses comparing battery trucks to electric road systems (ERS) highlight the complexity of infrastructure investment decisions, especially in cross-border settings [3, 10]. For instance, [3] emphasizes that without coordination, smaller countries may opt out of ERS investments, even if the overall EU welfare would benefit from collaboration, underscoring the need for integrated digital solutions that can bridge national and sectoral divides.
Cost-benefit analysis (CBA) offers a systematic framework for evaluating the economic, financial, and societal viability of complex integrated solutions like the one proposed herein. In European projects, standardized CBA methodologies have been developed to assess smart grid and sustainable energy projects. Notably, the European Commission Joint Research Centre (EC JRC) [11] outlines a multi-step CBA approach tailored for smart grid investments in developing countries, while [12] details a similar framework applied to a smart grid pilot in Isernia, Italy, focusing on distribution system automation. Both methodologies advocate a sequential analytical process, encompassing scope definition, benefit identification, monetization, and uncertainty assessment.
However, applying these CBA frameworks to comprehensively evaluate platforms that simultaneously manage energy and transport systems, particularly those leveraging advanced AI across multiple sectors and national borders, requires careful and detailed consideration of unique benefit streams and cost structures. While existing frameworks provide a robust foundation, their application must rigorously quantify the specific contributions of AI functionalities and cross-sectoral synergies. Even recent evaluations of integrated e-mobility systems often remain confined to national boundaries or siloed sectors [13, 14]. Studies assessing the economic viability of V2G strategies and smart charging [14], or real-time optimization for EVs [15, 16], often focus on partial benefits and neglect the compounded synergies across sectors and borders that a unified, AI-driven digital platform could unlock.
In this work, we propose a CBA framework to assess an AI-driven operational digital platform (ODP) that integrates EVs/ETs, RES, and the grid. While many platforms focus on specific aspects such as EV smart charging [17] or RES forecasting [18], the ODP holistically coordinates charging station availability, dynamic routing for EV/ET fleets considering energy constraints and battery health, demand response programs, and renewable energy forecasting within a unified, intelligent ecosystem. Specifically, this integrated management, orchestrated by the ODP’s AI engine is designed to boost grid flexibility by dynamically aligning EV/ET charging patterns with RES availability, improve overall system reliability through predictive load balancing and congestion management, and yield substantial economic benefits via optimized energy procurement, trading, operational cost reductions, and potentially extended EV/ET battery lifespans. To quantify these outcomes, we apply a seven-step CBA framework, thoroughly informed by [11] and [12]. Our application of this framework emphasizes the careful quantification of benefits directly attributable to the ODP’s capabilities (e.g., enhanced forecasting accuracy, superior optimization outcomes) and the synergistic effects of cross-sectoral integration, ensuring that AI-specific development, operational costs, and performance variabilities are explicitly addressed. This detailed application ensures alignment with EU best practices while providing a thorough assessment of this specific AI-driven, multi-sector and cross-border platform. We illustrate the framework to a comprehensive ODP deployments in Austria, Hungary, and Slovenia, comparing quantified benefits against investment and operational expenditures.
The main contributions of this paper are:
i)
Proposing a seven-step CBA framework to evaluate an AI-driven cross-sector ODP, explicitly identifying and monetizing AI-specific cost drivers (e.g., predictive analytics, optimization) and quantifying AI-driven benefit improvements (such as short-term RES forecasting and EV/ET charging optimization) as they propagate through the integrated transport–energy system.
ii)
Demonstrating a multi-layered, AI-driven ODP architecture by enumerating its core AI modules for RES forecasting, battery-health–aware EV charging, E-VRPTW routing for electric trucks, and co-optimization of grid ancillary services, and showing how verifiable performance metrics (e.g., improvement in price forecasting, reduction in battery degradation). This helps quantify economic, reliability, and environmental benefits, while addressing large-scale computational considerations (e.g., real-time neural network inference and parallelized mixed-integer programming).
iii)
Applying the proposed framework to real-world case studies in Austria, Hungary and Slovenia by utilizing country specific input data and projections (CAPEX, OPEX, RES capacity factors, EV/ET fleet sizes and their projections, and wholesale price time series) and computing present-value streams for each benefit (economic, reliability and environmental) alongside CAPEX/OPEX.
iv)
Deriving actionable policy and managerial insights (e.g., data-infrastructure requirements for 2026–2035, applicability to other EU Member States, and critical success factors such as interoperability standards and regulatory incentives for EV/ET aggregation) to inform Fit for 55 and REPowerEU objectives.
The remainder of the paper is structured as follows: Section 2 reviews related work. Section 3 presents the AI-driven ODP architecture, and the proposed CBA framework. Section 4 presents the CBA steps with a case study and application. Section 5 discusses the main results of the study and some of the limitations. Finally, Section 6 offers conclusions and future directions.
2 Related work
The pursuit of sustainable energy transitions has spurred significant research into the application of AI for optimizing the complex interplay between electric mobility, RES, and power grid operations [19, 1]. On the one hand, increasing reliance on digital platforms and smart charging programs can increase the risks of cyberattacks and natural disaster impacts on power systems [20]. The integration of EVs/ETS into the power system, which is overwhelmingly dominated by RES, has been the focus of several recent comprehensive reviews. Ref. [21] analyzes the economic dispatch and risk management challenges posed by large-scale plug-in hybrid EV/ET integration into power systems, emphasizing the need for coordinated scheduling strategies and robust optimization under uncertainty. Similarly, [22] provides a state-of-the-art review on the impact of fast EV charging on the utility sector, highlighting critical concerns such as load spikes, grid congestion, and the lack of regulatory readiness for widespread EV adoption. Flexibility needs in power systems with high levels of EV penetration and RES are reviewed by [23], who underscores the importance of adaptable system architecture to ensure reliability under dynamic conditions. On the other hand, recent literature underscores the transformative potential of AI in these domains, providing a robust academic foundation for the development and evaluation of integrated solutions like the AI-driven ODP proposed in this paper [24, 25, 26, 27]. These studies affirm that AI enhances the management of EVs, facilitates deeper RES integration, and enables more efficient and resilient grid operations, leading to considerable economic and environmental benefits. While the consensus points towards the positive impact of AI, the literature also acknowledges practical challenges such as data quality, system interoperability, cybersecurity, and the scalability of AI solutions, which are critical considerations addressed in the ODP’s design.
A foundational understanding of AI’s role in the broader energy transition is provided by Ref. [28], who conducts a comprehensive review highlighting AI’s contributions to RES deployment, energy efficiency enhancements, improved grid stability through smart grid integration, and the management of demand-side dynamics, including innovative vehicle-to-grid (V2G) systems. Their findings directly support the ODP’s core objective of orchestrating EVs, ETs, RES, and grid operations through sophisticated AI-driven forecasting and optimization. Complementing this, the survey by Ref. [29] on AI in the renewable energy domain identifies key AI-based strategies, such as neural networks and advanced optimization algorithms, that bolster the integration of diverse renewable sources like solar and wind. This broad perspective informs the selection and design of the ODP’s AI components, including its Long Short-Term Memory (LSTM) models for forecasting and mixed-integer linear programming (MILP) algorithms for optimization, aiming for robust and scalable solutions.
Focusing specifically on electric mobility, the systematic reviews by [30, 26, 31] on AI integration into EV energy management systems reveals significant advancements. Ref. [30] analysis of numerous studies highlights progress in energy optimization, intelligent route planning, charging demand forecasting, and predictive maintenance for EVs, primarily leveraging machine learning, deep learning, and genetic algorithms. On the distribution network level, Ref. [26] conducts an in-depth impact analysis of integrating projected EV charging loads into low-voltage systems, revealing significant concerns about voltage stability and transformer loading. Earlier work by Ref. [31] also provides a chronological review of EV/ET interactions with distribution systems, laying out both technical and regulatory gaps. The anecdotal analyses of AI techniques, by enhancing EV performance and battery health, directly underpin the ODP’s smart charging functionalities and logistics optimization modules, which notably incorporate considerations for battery degradation. The emphasis in their review on real-time adaptation to dynamic driving conditions and charging availability further informs the ODP’s AI-driven routing and intelligent charging strategies, aiming to maximize user benefits and overall system efficiency.
The challenge of integrating large-scale RES into existing energy systems is another critical focus area where AI demonstrates considerable promise. Ref. [32] investigates AI’s pivotal role in managing RES within multi-energy systems, particularly for achieving carbon neutrality. They emphasize AI’s capacity for intelligent operational control, achieved through precise forecasting of RES generation and effective demand-response strategies. This aligns closely with the ODP’s objective of synchronizing EV/ET charging demand with intermittent RES availability, thereby minimizing curtailment and maximizing the utilization of clean energy. The practical applications and identified limitations discussed by [32], such as the need for sophisticated and adaptive AI methodologies, provide valuable insights that have shaped the design of the ODP’s AI engine.
Parallel to advancements in RES integration and e-mobility, AI is recognized as a key enabler for the fundamental transformation of electrical grids into smarter, more resilient, and sustainable energy infrastructures. Ref. [33] explores this transformative impact, tracing the evolution from centralized to increasingly decentralized grid structures. Their work highlights AI’s crucial contributions to real-time data analytics, predictive maintenance of grid assets, and the optimization of demand-response programs. These AI-driven capabilities significantly enhance overall grid efficiency, operational flexibility, and resilience, directly corresponding to the ODP’s functions of improving grid stability and reliability through proactive, intelligent control strategies. The frameworks discussed for managing decentralized networks and ensuring interoperability inform the design of the ODP’s middleware layer and its approach to facilitating cross-border energy coordination. Such grid transformation efforts naturally lead to questions of economic justification.
Consequently, the economic and environmental viability of AI-driven smart energy systems becomes a crucial aspect for their practical implementation. Ref. [34], in its analysis of the InteGRIDy project, addresses this by conducting a detailed cost and environmental benefit assessment of integrating renewable energy and smart technologies. Their study quantifies tangible economic savings and significant environmental benefits, such as reductions in CO2 emissions. This work provides a pertinent methodological precedent and empirical support for the comprehensive seven-step CBA framework employed in our evaluation of the ODP, particularly in its approach to quantifying diverse benefits like reduced operational costs, enhanced market revenues, and monetized emission reductions across multiple national contexts.
Overall, this body of research underscores the significant potential of AI to revolutionize energy and transport systems. While individual studies demonstrate clear benefits in specific domains such as EV management, RES forecasting, or grid optimization, a clear opportunity remains for holistic platforms, like the proposed ODP, that integrate these functionalities to achieve synergistic, system-wide gains. The existing literature thus provides a strong impetus for the development and rigorous evaluation of such comprehensive AI-driven solutions. This study distinguishes itself by conducting a comprehensive CBA of a holistic, AI-driven ODP spanning multiple countries and sectors, with a particular focus on explicitly quantifying the incremental benefits derived from the integrated AI functionalities, an aspect often treated qualitatively or with a narrower scope in prior work. Our work on the CBA of ODP that integrates EVs/ETS, RES and grid responds to this by not only proposing such an integrated platform but also by conducting a detailed cost-benefit analysis to quantify its overall value proposition, directly addressing practical implementation challenges related to data management, AI model integration, and scalability, which are key considerations in our design.
3 AI-driven ODP and the CBA framework
This section presents the AI-driven ODP system architecture, an integrated platform for EV, ET, RES, and grid coordination, followed by the proposed CBA framework used to assess the financial and economic feasibility of the investment. Section 3.1 describes the multi-layered ODP architecture and AI engine, while Section 3.2 introduces the structured seven-step CBA framework.
3.1 ODP system architecture and AI engine
The AI-driven ODP is conceptualized as a multi-layered, interoperable platform designed to integrate a diverse array of technologies and stakeholders across the energy and transport domains. Its architecture, detailed in Table LABEL:tab_architecture, is structured into five principal layers:- Perception, Data, Middleware, Business, and Application. This layered design promotes modularity, scalability, and maintainability. The robustness of AI components is ensured through continuous monitoring, periodic retraining pipelines to counter data/concept drift, and leveraging scalable cloud infrastructure for computational needs, ensuring high availability and performance even with increasing data volumes and model complexity.
| Layer | Requirement | Type of technology / AI component | Main required technologies / AI model specification |
| Perception layer | Collect data on vehicles | EV/ET sensors | BMS sensors, GPS, OCPP-compliant charge monitors |
| Collect data on RES | RES sensors | Pyranometers, anemometers, power meters | |
| Collect data on environmental conditions | Environmental sensor arrays | IoT weather nodes, multi-sensor modules | |
| Collect data on weather conditions | Weather sensors | Integrated weather stations, satellite APIs | |
| Transmit data | Connectivity solutions | 5G, LTE-M/NB-IoT, LoRaWAN, Satellite | |
| Data layer | Receive data | Communication protocols | MQTT, HTTP/HTTPS, OPC-UA, AMQP |
| Communicate data | Message queues | Apache Kafka, RabbitMQ | |
| Store data | Databases | PostgreSQL, MongoDB, InfluxDB, TimescaleDB | |
| Store big data | Data lakes/warehouses | Azure Data Lake, Amazon S3, Hadoop HDFS | |
| Move data | Data integration, ETL | Apache NiFi, Talend, Azure Data Factory, AWS Glue | |
| Middleware layer | Integrate with existing systems | Integration middleware | FIWARE Context Broker, MuleSoft, Apache Camel |
| Provide clean APIs | API management | FIWARE Wilma/Keyrock, Apigee, AWS API Gateway | |
| Integrate with EU data spaces | Data space connectors | FIWARE Connector, IDS Connectors, Gaia-X | |
| Collect market data | Market integration middleware | EPEX SPOT API, Nord Pool API, ENTSO-E API | |
| Business layer (AI Engine Core) | Predict grid load, RES, prices | AI forecasting models | LSTM-RNNs, multi-head attention, NWP features, Azure ML, cloud GPU |
| Optimize EV/ET charging, V2X/V2B | AI optimization, RL agents | MILP/MIQP (CPLEX, Gurobi), DQN/PPO (RL), Python APIs, battery models | |
| Manage grid flexibility | Demand response management | OpenADR 2.0b tools, DR event prediction, load aggregation algorithms | |
| Optimize EV/ET routing | Route optimization engines | Electric VRP solvers, Google OR-Tools, metaheuristics, battery-aware routing | |
| Weather forecast integration | Advanced weather modeling | NWP models (ECMWF, GFS), AI correction (GBM, NN), sensor fusion | |
| Application layer | User interaction | Mobile/web apps | React Native, Flutter, charge planning, fleet dashboards |
| Navigation to charging points | Navigation systems | Google Maps API, Mapbox, HERE, ODP charger data integration | |
| ODP insights | Analytics/visualization | Power BI, Grafana, Tableau, KPI dashboards | |
| Administration tools | Management dashboards | Grafana, Kibana, model drift detection, admin UI |
1.
Note: The ODP architecture is based on the EU-funded BEGONIA project; all layers are described in detail in Deliverable D3.1 [35].
The Perception layer is responsible for real-time data acquisition from a multitude of distributed assets, including EVs, ETs, RES installations (solar PV, wind turbines), charging infrastructure, and environmental sensors. This layer employs various sensors (BMS data, GPS, smart meters, pyranometers, anemometers) and communication technologies (5G, LTE-M, LoRaWAN) to ensure robust and timely data transmission. Data is collected at granular resolutions (e.g., 1-minute intervals for critical parameters). The Data layer manages the ingestion, processing, storage, and secure handling of these voluminous and heterogeneous data streams. It leverages communication protocols (MQTT, OPC-UA), message queuing systems (Apache Kafka), and a hybrid data storage architecture comprising time-series databases (InfluxDB, TimescaleDB) for operational data and data lakes (e.g., Azure Data Lake) for large-scale analytics and AI model training. Data quality validation, anomaly detection, and missing data imputation are integrated processes within this layer to ensure robust inputs for the AI engine.
The Middleware layer provides crucial interoperability services, enabling seamless integration with existing utility systems (SCADA/DMS/EMS), energy market platforms (e.g., EPEX SPOT, Nord Pool APIs), and emerging EU data spaces (e.g., Gaia-X, IDS connectors). It utilizes enterprise integration patterns, API management gateways, and connectors designed for specific data exchange standards (e.g., ENTSO-E data models, Open Charging Protocols). This layer ensures secure, standardized, and real-time data flow necessary for cross-sectoral coordination.
The Business layer constitutes the AI engine core of the ODP, housing the advanced algorithms and models that drive intelligent decision-making. Key AI components include:
Forecasting models: Long Short-Term Memory (LSTM) based Recurrent Neural Networks (RNNs), enhanced with multi-head attention mechanisms, are employed for forecasting EV/ET charging demand, RES generation, and grid load at various temporal and spatial resolutions (e.g., 15-minute intervals up to 72 hours ahead). These models capture complex temporal dependencies and non-linear relationships inherent in energy data, using inputs like historical load/generation, high-resolution weather forecasts (temperature, irradiance, wind speed, cloud cover from numerical weather prediction models and local sensors), calendar effects (day-of-week, holidays), and real-time market signals. Model training is performed on extensive historical datasets (e.g., 2-3 years of data from ENTSO-E, national TSOs/DSOs, and representative EV/ET usage profiles) using scalable cloud-based GPU resources. Performance is continuously validated through rolling-window cross-validation and backtesting against established baseline methods (e.g., ARIMA, persistence), achieving a target Mean Absolute Percentage Error (MAPE) below 5% for day-ahead RES and load forecasts, demonstrating a significant uplift in predictive power essential for subsequent optimization steps. Energy price forecasting models, potentially hybrid approaches combining statistical models (e.g., ARIMA-GARCH) with machine learning for residual analysis, are also developed to predict day-ahead and intra-day electricity prices.
EV/ET smart charging and V2X dispatch: mixed-integer linear programming (MILP) or mixed-integer quadratic programming (MIQP) formulations optimize charging/discharging schedules for individual vehicles and aggregated fleets. The objective functions are multi-faceted: minimize total energy cost for users, maximize self-consumption of local RES, minimize carbon emissions, reduce peak grid load, and critically, promote battery longevity. Battery degradation is explicitly incorporated into the objective by penalizing charging strategies known to accelerate degradation(e.g., frequent deep discharges, high state-of-charge (SoC) extremes, high C-rates). For instance, additional cost term is included, where represents a battery aging model based on SoC range, C-rates, and temperature, informed by established research [36, 37, 38]. The AI optimizer thus actively minimizes this degradation cost as part of its multi-objective optimization. Constraints include battery SoC limits (), user-defined departure times and required SoC targets, maximum charging/discharging rates, local grid capacity limits, charger availability (from Perception layer), and electricity tariff structures. Commercial solvers like CPLEX or Gurobi are employed via Python APIs. For very large problem instances (thousands of vehicles), decomposition techniques (e.g., column generation) or distributed optimization methods ensure solutions within operational timeframes (minutes for day-ahead scheduling, seconds for real-time adjustments). In addition, V2G/V2B services are managed by reinforcement learning (RL) agents (e.g., Deep Q-Networks or Proximal Policy Optimization). These agents are trained on high-fidelity simulations of real-time energy markets (reflecting bidding rules, settlement periods for frequency regulation services) and grid conditions. The RL reward function incentivizes revenue from ancillary services while safeguarding battery health and user mobility needs. Simulations incorporate historical market data, realistic network constraints, and stochastic elements to approximate real-world uncertainty.
ET route and logistics optimization: Sophisticated vehicle routing problem (VRP) algorithms, specifically tailored for electric fleets (E-VRPTW – Electric VRP with time windows and charging stops), are implemented. These algorithms optimize routes to minimize total operational cost (combining energy consumption, time, and potential battery degradation costs) while considering real-time traffic conditions, terrain (for energy consumption estimation), payload, dynamic weather conditions, charging station availability, charging duration, and potential queuing times at charging points. The integration with battery degradation models ensures that routing decisions do not inadvertently lead to accelerated battery wear. Algorithms are based on advanced metaheuristics (e.g., tabu search, genetic algorithms) or commercial solvers (e.g., Google OR-Tools). The performance is evaluated by metrics such as route efficiency (km/kWh), total travel time, and number of required charging stops, compared to static or less intelligent routing systems.
Demand response (DR) management: AI-driven tools automate the identification and aggregation of flexible EV/ET loads, predict their response potential based on historical patterns and current vehicle states, and manage their participation in DR events or ancillary service markets. This adheres to standards like OpenADR 2.0b. The system uses AI-driven baseline calculation methods to accurately measure DR event effectiveness and optimize dispatch signals to individual vehicles or aggregated charging stations.
These AI models are developed using frameworks like TensorFlow, Keras, and PyTorch, integrating with optimization solvers such as CPLEX or Gurobi via Python APIs. Model training utilizes high-quality historical and real-time data from the Data layer. Crucially, the ODP incorporates continuous operational monitoring, performance validation metrics (e.g., forecast accuracy, optimization solution quality, solver runtime), and scheduled retraining pipelines to proactively address potential data drift, concept drift, or model degradation over time, ensuring sustained AI model efficacy and robust decision-making. Cloud-native scalable infrastructure supports the computational demands of large-scale AI model training and real-time inference.
Finally, the Application layer delivers user-facing interfaces and services. This includes intuitive mobile and web applications for EV/ET drivers and fleet managers (e.g., charge planning, station booking, real-time route guidance with charging recommendations), as well as comprehensive analytics dashboards and administrative consoles for system operators, charging point operators (CPOs), and energy providers. These dashboards provide transparent insights into system performance, AI model outputs, and optimization outcomes. This layered, AI-centric architecture allows the ODP to ingest vast quantities of real-time data, transform it into actionable intelligence through its AI engine, and orchestrate coordinated actions across the transport and energy sectors. The platform’s design inherently facilitates interoperability, dynamically aligning EV/ET charging with RES availability and grid conditions, while integrating market signals to optimize economic, environmental, and battery health outcomes, thereby creating an intelligent nexus for sustainable energy and mobility.
3.2 Cost-benefit analysis framework
To evaluate the proposed AI-driven ODP, we employ a structured seven-step CBA framework, guided by European Commission smart grid guidelines [11] and lessons from pilot studies such as Isernia [12]. This framework systematically captures AI-related costs (platform development, sensors and IoT devices, model training and maintenance) and the benefits directly attributable to AI-driven performance (e.g., improved forecast accuracy, optimization savings). The framework starts with a precise definition of the project, its objectives, engineering features, geographic scope, stakeholders, and expected socio-economic impact, and delineation of its functional scope. Next, it maps the enabling technologies (especially AI modules) to the key functions they perform, and links those functions to the spectrum of benefits they generate. Subsequently, each category of benefit is quantified and monetized, and all associated costs (capital and operating) are tallied over the project’s life cycle. Finally, total benefits and costs are compared, and a rigorous sensitivity analysis, including scenario variations and Monte Carlo simulation, is conducted to assess how results vary under key uncertainties (such as AI performance and adoption rates). Table 2 summarizes these steps.
| Step | Procedure |
| 1 | Define the project (scope, objectives, engineering features, AI functionalities, and their specific role). |
| 2 | Map technologies (including AI modules and their specific functionalities) to platform functions. |
| 3 | Map platform functions to economic, reliability, and environmental benefits. |
| 4 | Monetize and quantify benefits (linking ODP performance to benefit magnitude, with transparent assumptions). |
| 5 | Quantify costs (CAPEX and OPEX, including detailed AI system development, training, and maintenance costs). |
| 6 | Compare costs and benefits (NPV, BCR). |
| 7 | Conduct sensitivity analysis and risk assessment (including discount rate, technology adoption variability parameters, and Monte Carlo simulation). |
With this framework, Step 1 establishes the project’s scope, objectives, geographic coverage, stakeholders, and counterfactual baseline. Step 2 links each ODP technology layer and AI component to its operational function (data acquisition, integration, forecasting & optimization, decision support). Step 3 identifies how these functions generate economic, reliability, and environmental benefits. In Step 4, each benefit is converted into monetary terms over the project horizon using transparent formulas and country-specific data. Step 5 aggregates all capital and operating expenditures necessary for deployment and operation, including AI model development, sensors, cloud services, and personnel. Step 6 computes net present value (NPV) and benefit–cost ratio (BCR) by discounting all cash flows, reflecting societal time preferences. Finally, Step 7 tests the robustness of the results through scenario variations and Monte Carlo simulation, and identifies key risks with corresponding mitigation strategies. Section 4 applies these steps in detail to the Austria–Hungary–Slovenia case study.
4 Cost-benefit analysis steps with case study
We now apply the seven-step CBA framework to the AI-driven ODP project, evaluating its deployment across Austria (AT), Hungary (HU), and Slovenia (SI). For the economic appraisal, all future cash flows are expressed in constant 2025 Euros and discounted to present value using a social discount rate of 4%. This choice is consistent with EU recommendations for long-term public infrastructure projects [39]. The results highlight the value created by the ODP’s AI-driven features and provide insight into the project’s economic viability under various scenarios.
Step 1: Define the project
The project under consideration is the deployment and operation of the AI-driven ODP for the integrated management of EVs, ETs, RES, and the electric grid across three countries: Austria, Hungary, and Slovenia. The primary objective of this project is to leverage real-time data and advanced AI (for forecasting and control) to enhance grid flexibility, reduce energy procurement and operational costs, minimize greenhouse gas emissions and local air pollutants, improve power system reliability, and enable seamless interoperability between transport electrification and power system operations. The implementation scale is national/regional within these countries, which have been chosen for their diverse EV/ET adoption rates, renewable energy capacities, and grid characteristics, thereby providing a representative range of EU contexts for analysis.
Key engineering features of the ODP critical to the project include: (a) AI-powered predictive algorithms for demand forecasting (EV/ET charging, overall grid load), RES generation forecasting, and energy price forecasting. These provide high accuracy (e.g., day-ahead load and RES forecasts with MAPE ), superior to conventional methods, directly translating to improved operational decisions and cost savings. (b) Optimization engines (MILP-based solvers and RL agents) for smart charging scheduling, V2X services, electric fleet logistics (routing and charging), and demand response dispatch. These engines optimize multi-objective functions, including cost reduction, emissions reduction, and battery health preservation, thus enabling coordinated control strategies that balance the grid and maximize economic returns. (c) IoT infrastructure comprising a dense network of sensors and smart meters to provide granular, real-time monitoring of vehicles(battery SoC, location, charging status), charging stations (availability, power flow), grid nodes (voltage, load), and RES installations (output, status). This high-resolution data collection is foundational for the AI engine’s situational awareness. (d) High-speed, reliable communication networks (5G, LTE-M) ensuring low-latency data transfer and control signal dissemination. This is critical for real-time AI inference and control, allowing the ODP to rapidly respond to changing conditions (such as sudden RES output fluctuations or grid contingencies). (e) Robust middleware integration with existing energy management systems (utility SCADA/DMS/EMS) and market platforms. By using standardized interfaces and data exchange protocols, the ODP can seamlessly interoperate with legacy systems and participate in energy markets (for energy trading and ancillary services), thus embedding the platform within the broader energy ecosystem.
The primary stakeholders expected to benefit from this ODP include: EV owners and ET fleet operators (through optimized charging costs, reduced range anxiety, improved charging availability, extended battery lifespan, and efficient routing); charging point operators (CPOs) (via enhanced asset utilization and new service offerings); distribution and transmission system operators (DSOs/TSOs) (through reduced grid congestion, deferred network reinforcement investments, improved grid stability, and enhanced situational awareness due to AI-driven insights); renewable energy producers (via reduced curtailment and better market integration of variable generation); and society at large (through reduced carbon emissions, improved air quality, and enhanced energy security).
The baseline scenario (counterfactual case without the ODP) assumes that EV and ET adoption and renewable deployment continue to grow in the three countries, but in a conventional manner: EV charging remains largely uncoordinated (users charge at will, leading to sharper peaks and low overnight utilization), V2G and other vehicle-to-grid services have minimal participation, RES curtailment issues persist at times of excess generation, and grid operators have limited visibility/control over distributed EV chargers and other flexible resources. Essentially, the baseline reflects current practices where the transport and energy sectors operate with only rudimentary interaction and rely on existing, less sophisticated forecasting and control tools. The cost-benefit analysis will evaluate the incremental impacts of implementing the ODP relative to this baseline, thereby isolating the net benefits attributable to the platform’s advanced AI-driven coordination.
Step 2: Map technologies to functions
In this step, each core technology component of the ODP, particularly its AI modules as detailed in Section 3.1, is systematically mapped to the specific operational functions it enables. The ODP integrates a comprehensive suite of technologies (sensors, communication networks, data platforms, AI algorithms, user interfaces), which together deliver several key functions for the system’s operation:
(A) Advanced data and information collection & processing: Real-time gathering, validation, and processing of granular data from EVs/ETs (e.g., battery SoC, location, charging patterns), RES installations (generation output, availability), grid sensors (voltage levels, loading), charging infrastructure status, and external sources (weather, market market prices). This function is enabled by IoT sensors, connectivity solutions, and the ODP’s Data layer; it ensures that the AI engine has accurate and timely information about the system’s state at all times.
(B) Enhanced interoperability and system integration: Seamless data exchange and coordinated operation among diverse systems and stakeholders (vehicles, grid operators, CPOs, energy markets, end-users). This is facilitated by the Middleware layer and use of standardized communication protocols and interfaces (e.g., OCPP for charging, OpenADR for demand response signals, API gateways for market data). Enhanced interoperability ensures that decisions made by the AI core can be effectively communicated and executed across different subsystems in real time.
(C) AI-driven forecasting, optimization, and control: This core function is where the ODP’s AI engine directly drives value. It has two sub-functions:
i) Intelligent forecasting: The AI models (LSTMs, ML models, etc., as described in Section 3.1) provide highly accurate predictions for EV/ET charging demand, RES generation, grid load, and electricity prices. These forecasts improve upon traditional methods, reducing uncertainty and prediction error in the system’s operation.
ii) Smart load management and grid balancing: AI optimization algorithms (MILP, RL, E-VRPTW, as detailed in Section 3.1) enable dynamic smart charging/discharging of EVs/ETs, optimal dispatch of V2X services for grid support, dynamic EV/ET routing considering energy and battery health, and coordinated demand response actions. These control decisions continuously balance supply and demand, flatten load curves, alleviate grid constraints, and maximize the use of renewable energy.
(D) Decision support and automation: User-friendly interfaces (mobile apps, dashboards) provide actionable insights and recommendations to end-users (e.g., optimal charging times) and system operators (e.g., congestion alerts and recommended mitigation actions). Automated control systems execute optimized schedules and dispatch signals directly generated by the AI engine, ensuring real-time response and reducing manual intervention.
Table 3 provides a high-level mapping of the ODP’s major technological assets and AI components to these functional categories A–D. For instance, IoT sensors and connectivity primarily support data collection (Function A). The integration middleware and communication protocols enable Function B. The AI forecasting models and optimization algorithms directly drive Function C (intelligent forecasting and control), while the user interfaces and automated control modules deliver Function D (decision support and action). This detailed mapping ensures that all critical functions necessary for the ODP’s success are supported by specific, well-defined technological capabilities, with the core AI engine acting as the central intelligence orchestrating these functions.
| Technology (asset) / AI component | A | B | C | D |
| IoT sensors (EV, ET, RES, weather, grid) | ✓ | |||
| Edge gateways | ✓ | ✓ | ||
| 5G / LTE-M / LPWAN connectivity | ✓ | ✓ | ||
| Message queue systems (Kafka) | ✓ | ✓ | ||
| Communication protocols (MQTT, OCPP, OpenADR) | ✓ | ✓ | ||
| Microservice architecture | ✓ | ✓ | ||
| Data integration middleware (FIWARE) | ✓ | |||
| Data lake & time-series databases | ✓ | ✓ | ||
| AI forecasting models (LSTM, etc.) | ✓ | ✓ | ||
| AI Optimization algorithms (MILP, RL, E-VRPTW) | ✓ | ✓ | ||
| Demand response management tools | ✓ | ✓ | ✓ | |
| External APIs (market, grid, weather) | ✓ | ✓ | ✓ | |
| User interfaces/dashboards (Mobile Apps, Web) | ✓ | |||
| Automated control modules | ✓ | ✓ |
This systematic mapping confirms that the ODP’s technology stack, with its integrated AI engine, comprehensively addresses the functional requirements for data acquisition, system integration, intelligent analysis and control, and user interaction. This forms a robust foundation for optimizing energy and mobility services, creating synergistic benefits by enabling real-time, coordinated actions across the energy and transport domains that are not possible with siloed systems.
Step 3: Map functions to benefits
This step identifies the specific categories of benefits generated by each core ODP function. The anticipated benefits are categorized into three primary domains:
i) Economic benefits: Direct financial gains or cost savings, encompassing reduced operational expenditures for grid operators and end-users, new revenue streams from energy services (e.g., ancillary services, energy arbitrage), enhanced energy efficiency, and optimized asset utilization.
ii) Reliability benefits: Improvements in power system reliability and resilience, leading to reduced outage costs for consumers and businesses, enhanced service quality for end-users, and potentially deferred or optimized investments in grid infrastructure due to better load management.
iii) Environmental benefits: Reductions in greenhouse gas (GHG) emissions (primarily CO2) and local air pollutants (SOx, NOx, PM10), contributing to climate change mitigation, improved public health, and greater integration of clean renewable energy.
Each ODP function (A-D) contributes to one or more of these benefit categories. For example, Advanced data collection and processing (A) provides the necessary real-time data for precise grid state estimation and load forecasting, enabling more efficient asset utilization and informed decision-making, which directly translates to economic efficiency and enhanced system reliability. Enhanced interoperability (B) allows for dynamic coordination between different market participants and grid entities, enabling new revenue streams from flexible assets and reducing system imbalances. The core AI-driven forecasting, optimization, and control functions (C) are pivotal in generating significant economic benefits (e.g., reduced energy procurement costs through peak load shaving and optimized energy trading facilitated by highly accurate price predictions, and increased revenues from ancillary service provision). These AI functions also yield substantial reliability benefits (e.g., prevention of grid overloads, improved voltage/frequency stability through predictive control, and reduced RES curtailment by matching demand with variable generation). Furthermore, these AI functions directly contribute to environmental benefits by maximizing the use of RES, optimizing EV charging for lower emission periods, and reducing fossil fuel-based generation for balancing. Decision support and automation (D) translates AI-generated insights into tangible outcomes, yielding direct economic benefits for users (e.g., lower charging costs due to optimized schedules, time savings from efficient routing) and operators (e.g., automated response to grid events), and indirectly fostering environmental benefits by promoting behaviors aligned with RES availability and lower carbon intensity.
Table 4 qualitatively summarizes how the ODP’s key functions map to specific benefit mechanisms.
| Function | Economic | Reliability | Environmental |
| A. Advanced data & info. collection/processing | ✓ | ✓ | (i) |
| B. Enhanced interoperability/system integration | ✓ | ✓ | ✓ |
| C. AI forecasting, optimization & control | ✓ | ✓ | ✓ |
| D. Decision support & automation | ✓ | (i) | ✓ |
This mapping highlights that economic benefits are pervasively driven by all functional areas, with particularly strong and direct contributions from the AI engine (Function C) and effective user/system engagement through transparent decision support (Function D). Reliability benefits primarily stem from improved situational awareness (Function A), seamless integration enabling coordinated responses (Function B), and intelligent grid management capabilities (Function C) that anticipate and mitigate grid issues. Environmental benefits accrue mainly from functions that facilitate higher RES penetration and smarter, cleaner energy consumption patterns for EVs/ETs (Functions B, C, and D), minimizing reliance on fossil fuels.
Step 4: Monetize benefits
Having identified the benefits qualitatively, this step involves quantifying each benefit in monetary terms over the analysis horizon (2026–2035). Specific metrics, mathematical formulations, and assumptions are established for each benefit category. The details are presented in A.1, based upon standard valuation methodologies, relevant literature [11, 40], and country-specific data. The quantification of AI’s contribution is a central focus, linking specific AI model performance metrics and optimization efficiencies to the magnitude of the derived benefits. For each relevant benefit, the improvement margin over the baseline is directly tied to the enhanced capabilities, either superior forecasting accuracy or more efficient operational optimization, provided by the specific AI components of the ODP detailed in Section 3.1. To demonstrate the ODP’s impact in varied national contexts, the proposed CBA framework is applied to Austria (AT), Hungary (HU), and Slovenia (SI). These countries offer a compelling mix of varying grid structures, renewable energy penetration levels, and electric mobility adoption trajectories, providing a representative testbed for the ODP’s cross-border and multi-sectoral capabilities.
| Parameter | Austria (AT) | Hungary (HU) | Slovenia (SI) | Source/Justification |
| EV stock (2026, thousands) | 343.8 | 100.8 | 50.0 | National Registries, [41] |
| Annual EV Growth rate (CAGR 2026-35) | NECPs, [42] | |||
| ET stock (2026, thousands) | 4.4 | 0.4 | 0.3 | Pilot data, [43] |
| Annual ET growth rate (CAGR 2026-35) | Industry projections, [44] | |||
| RES capacity (Solar + Wind, 2026 GW) | 10.5 | 8.0 | 1.5 | National TSO data, [45] |
| Annual RES addition (MW/yr avg.) | 664 | 2750 | 150 | NECP targets, [46] |
| Avg. electricity retail price (EV, €/kWh) | 0.22 | 0.15 | 0.18 | [47], adj. for EV tariffs |
| Grid CO2 intensity (2026, g/kWh) | 120 | 280 | 190 | [48], projected |
| Value of lost load (VoLL, €/kWh) | 10.0 | 7.5 | 8.0 | Nat. regulator studies, [49] |
| Battery life Ext. (smart charging) | reduction in degradation rate | [37, 38, 36] | ||
| Social cost of carbon (€/tCO2) | Starts at €85 in 2026, rising to €120 by 2035 | EU guidance [50], linear interpolation. | ||
Table 5 summarizes key country-specific input assumptions for the baseline year 2026 and projected growth rates for the analysis period (details are presented in B). These assumptions are derived from recent national energy and climate plans (NECPs), statistical office data [51], reports from international agencies like the IEA [52], and specialized consultancies (e.g., BloombergNEF, McKinsey) to ensure their credibility and relevance to the respective national contexts. For instance, the EV/ET growth rates reflect ambitious but achievable national targets and industry projections, while RES capacity additions align with NECP goals. The assumed reduction in battery degradation rate due to ODP-enabled smart charging, which explicitly incorporates battery health objectives in its AI optimization algorithms (see Section 3.1), is monetized through calculating the net present value of deferred battery replacement costs or the increased residual value of the vehicle. This benefit, derived from the ODP’s AI-driven smart charging (Section 3.1) and validated by internal simulations against established battery aging models [36, 37, 38], involves managing charge/discharge cycles, temperature, and depth of discharge to minimize stress on the battery. Key benefit categories and their monetization approaches are summarized below. Detailed calculations of each benefit and the specific data required for these calculations are presented in Appendices from A.1.1 to A.1.8.
i) Reduction in operational downtime (ROD) cost: This is from avoided power outages or avoided unserved energy due to improved grid reliability and avoided EV/ET service disruptions. A conservative reduction in baseline outage energy could be avoided due to ODP systems enhancement (1) and AI-driven predictive analytics enhancement111For all economic, reliability, and environmental benefits listed here, readers are referred to Section 3.1 for a description of the AI-driven models and the architecture of the ODP that integrates EV, ET, RES, and the grid using different advanced AI layers.
enabling proactive grid management and faster fault restoration. This analysis could be benchmarked against improvements reported in analogous smart grid deployments [53].
ii) Revenue from optimized energy trading and ancillary services (ROETAS): This is an increase in revenue or cost savings from optimized participation in wholesale energy markets and ancillary services. This benefit is directly influenced by the forecasting accuracy of ODP’s AI models (2).
iii) Cost savings from demand response and peak load reduction (CSDR-PLR): This is the avoided costs from reducing system peak demand through coordinated EV/ET smart charging and DR programs orchestrated by the ODP’s AI. Peak load reductions of (3) are achieved by the ODP’s MILP-based optimization engine.
iv) Avoided energy curtailment (AEC): It is a monetized value of renewable energy utilized rather than curtailed due to the ODP’s intelligent management. A reduction in baseline curtailment (5) is projected based on simulations where the ODP’s AI (RES forecasting and demand-flexibility matching capabilities) dynamically aligns EV/ET charging with RES availability.
v) Fleet energy and operational savings (FES): This is a reduced energy consumption and operational costs for EV/ET fleets. This includes savings from AI-optimized routing (E-VRPTW algorithms in ODP, yielding an estimated reduction in energy/distance for relevant trips, based on pilot study analogs of AI-optimized fleet routing). A significant component is the monetized value of extended battery lifespan (estimated at reduction in degradation rate). This benefit is directly achieved by the ODP’s AI-driven smart charging algorithms, which explicitly incorporate battery health objectives (e.g., optimized SoC profiles, C-rate management, thermal pre-conditioning (6)) informed by established battery aging models [36, 37, 38].
vi) Grid stability and management cost savings (GSMS): This benefit is considered as reduced costs for DSOs/TSOs associated with grid regulation, voltage support, and congestion management. The benefit is calculated as grid stability savings and grid stability from forecasting ability of the AI-driven ODP (7). This estimation is based on the ODP’s AI-driven enhanced forecasting and proactive control capabilities, leading to operational efficiency gains analogous to those reported in advanced smart grid operational studies [53, 54] and analyses of AI-driven grid management platforms.
vii) CO2 emission reduction: This is a monetized benefit of greenhouse gas emission reductions attributable to the ODP’s operations, such as increased RES utilization, optimized EV/ET charging due ODP-enabled RES, EV, ET and awareness creation on the use of operational platform for social cost of carbon reduction (8).
viii) Reduction in air pollutant (RAP): This is a monetized health and environmental improvement from reduced local air pollutant emissions (9).
The detailed mathematical formulations and parameter derivations for the above benefits are provided in A.1. By summing the discounted annual benefits across the 10-year analysis period, we obtain the present value for each benefit category and the total benefits.
Applying the monetization formulas with country-specific assumptions yields the projected benefits for each nation.
Table 6 presents the total present value of benefits over , aggregated across the three countries and categorized by type. Detailed annual benefit projections, underpinning these present values, are also provided in B.
| Benefit category | Component | Present value (€ million) | Share of total benefits (%) |
| Economic benefits | Reduction in operational downtime (ROD) | 70.63 | |
| Optimized energy trading & ancillary services (ROETAS) | 148.20 | ||
| Demand response & peak load reduction (CSDR-PLR) | 170.80 | ||
| Fleet energy & operational savings (FES) | 75.61 | ||
| Avoided energy curtailment (AEC) | 292.99 | ||
| Subtotal economic benefits | 758.23 | 61.5 | |
| Reliability benefitsa | Grid stability & management savings (GSMS)b | 347.99 | |
| Subtotal reliability benefits | 347.99 | 28.2 | |
| Environmental benefits | CO2 emission reduction benefits | 89.07 | |
| Reduction in air pollutant benefits (RAP) | 38.64 | ||
| Subtotal environmental benefits | 127.71 | 10.4 | |
| Total monetized benefits (PV) | 1233.9 | 100.0 |
a
Reliability benefits are substantial. GSMS captures direct grid operational savings. ROD, while categorized under economic, also reflects improved reliability. AEC enhances stability by better integrating variable RES.
b
GSMS: Reduced costs for grid regulation, balancing services (not otherwise captured in ROETAS), and congestion management due to ODP’s predictive control and enhanced visibility.
Economic benefits constitute the largest share ( of total benefits), primarily driven by the AI-enabled optimized energy trading, demand response, and fleet operational savings. Reliability benefits, quantified mainly through grid stability and management savings (GSMS) at ), presents a significant portion of the benefits, indicating the ODP’s potential to enhance grid resilience and potentially defer traditional grid infrastructure investments. Environmental benefits also show a significant () benefit contribution, showing the ODP’s role in accelerating decarbonization.
Figure 1 illustrates the composition of total benefits, clearly depicting the dominant economic contributions and the significant environmental and reliability gains.
Country-specific analyses (based on the annual data presented in B and Monte Carlo simulation results in A.4) reveal nuanced variations in benefit distribution, reflecting national characteristics. For instance, Austria, with its higher initial EV adoption and ambitious electrification targets, sees proportionally greater benefits from fleet energy and operational savings (FES) and reduction in operational downtime (ROD). Hungary, with its significant RES expansion plans, benefits more substantially from avoided energy curtailment (AEC) and revenue from optimized energy trading and ancillary services (ROETAS) due to the greater flexibility needs. Slovenia, given its smaller scale but gains particularly from optimized ET logistics and enhanced cross-border grid balancing, which the ODP is designed to facilitate. These variations underscore the ODP’s adaptable value proposition across diverse EU member states.
Step 5: Quantify costs
This step involves a comprehensive enumeration and quantification of all costs associated with the implementation and operation of the ODP over the 2026–2035 period. Costs are categorized into Capital expenditures (CAPEX) and Operational expenditures (OPEX). Detailed breakdowns, differentiating between initial platform development and ongoing integration costs, are provided in A.2 (CAPEX) and A.3 (OPEX). Below, we explain their main elements:
CAPEX: These are primarily upfront investments incurred during the initial deployment phase (2026) and for system expansion in subsequent years. Major CAPEX components include hardware procurement (sensors, servers, communication devices), core software development for the ODP platform, and crucially, the initial development and training of the AI engine’s specialized models (e.g., LSTM architectures, MILP/RL solver integration, data pipelines for AI). System integration and project management costs are also included. A significant portion of CAPEX is incremental, related to the integration of new EVs, ETs, and RES installations annually into the ODP-managed ecosystem. Our estimates indicate a total discounted CAPEX of €561.3 million over the 10-year period, with approximately of the initial platform CAPEX directly attributable to AI model development and associated infrastructure (e.g., specialized GPU clusters for training).
OPEX: These are recurring annual costs necessary to operate and maintain the ODP. Key OPEX categories include cloud computing resources for AI inference and data processing, ongoing software maintenance (including continuous AI model monitoring, re-training pipelines to counter data/concept drift, and performance optimization), communication fees, robust cybersecurity operations, labor costs for technical staff (e.g., data scientists, AI engineers, system administrators), and hardware maintenance. Annual OPEX is projected to start at approximately €31.6 million (discounted) in the first year (2026), with an estimated of this directly related to AI operations and maintenance. The total discounted OPEX over 10 years is estimated at €315.9 million.
Figure 2 illustrates the indicative distribution of OPEX in the first year of operation (2026), demonstrating the allocation across various operational aspects, including those supporting AI functionality.
The total discounted investment cost (CAPEX + OPEX) for the ODP deployment and operation over the decade is approximately €877.2 million. Cost estimates are benchmarked against similar smart grid and digital platform deployments in the energy sector [12, 55] and include reasonable contingencies for unforeseen expenditures. The specific breakdown for AI-related costs captures both initial development (CAPEX) and ongoing operational and maintenance expenses (OPEX) essential for sustaining AI model performance in a dynamic environment.
Total costs (CAPEX and OPEX) for the aggregated three-country deployment amount to €877.2 million. CAPEX for shared platform development and core infrastructure, including AI engine development, was allocated to each country based on a composite weighting factor considering projected deployment scale (e.g., number of managed EVs/ETs, RES capacity integrated, and overall energy consumption), as detailed in A.2 and A.3. As a result, Austria accounts for approximately of total costs, Hungary , and Slovenia . Detailed cost breakdowns, including specific required items for AI system development and ongoing AI model maintenance, are comprehensively presented in A.2 and A.3, with annual cost projections provided in B.
Step 6: Compare costs and benefits
In this step, the total present value of monetized benefits is compared against the total present value of costs. Key financial performance indicators, including NPV and BCR, are calculated to assess the overall economic viability of the ODP investment. Table 7 summarizes the aggregated financial outcomes over the 10-year analysis horizon (2026–2035), discounted at .
| Category | Present value (€ million) |
| Total benefits | |
| CAPEX | |
| OPEX | |
| Total costs (CAPEX + OPEX) | |
| NPV (Total benefits Total costs) | 356.7 |
| BCR (Total benefits / Total costs) | 1.41 |
The analysis reveals that total discounted benefits (€1,233.9 million) significantly exceed total discounted costs (€877.2 million). This results in a positive NPV of approximately €356.7 million and a BCR of 1.41. A BCR greater than 1 indicates that the project’s benefits outweigh its costs, with every Euro invested generating €1.41 in benefits. This suggests that the ODP is a financially sound and economically justifiable investment, particularly considering the long-term societal and environmental returns.
Figure 3 provides a visual comparison of the aggregated costs and benefits by their primary categories. Economic benefits constitute the largest share, followed by substantial contributions from reliability and environmental benefits.
The positive NPV and BCR 1 strongly support the economic and financial feasibility for the ODP deployment. These results are based on conservative assumptions for benefit realization and cost estimations, detailed in A.2 and A.3, and include the explicit quantification of AI’s performance-driven contributions.
Step 7: Sensitivity analysis and risk assessment
To assess the robustness of the CBA results to inherent uncertainties in key input parameters, a comprehensive sensitivity and risk analysis was performed. This involved both scenario-based sensitivity testing and probabilistic risk analysis using Monte Carlo simulation. First, scenario-based analyses were conducted by varying critical benefit and cost parameters. Table 8 presents the impact of these variations on the project’s NPV and BCR. Crucially, this includes a specific scenario for AI adoption variability, where the effectiveness factors are reduced.
| Scenario | NPV (€ million) | BCR |
| Base Case | ||
| Total benefits + | ||
| Total benefits - | ||
| Total costs + | ||
| Total costs - | ||
| Discount rate (instead of ) | ||
| Discount rate (instead of ) |
The scenario analysis indicates that the ODP investment remains viable (BCR 1, NPV 0) even under reasonably pessimistic assumptions. The sensitivity analysis demonstrates a comfortable margin of safety for the investment. A Monte Carlo simulation ( iterations) was also performed. Key parameters such as EV/ET adoption rates, energy price volatility, RES generation variability, CAPEX/OPEX unit costs, AI model performance factors (e.g., ranges for forecast accuracy benefits, optimization efficiency gains), and the social cost of carbon were modeled as random variables with appropriate probability distributions (e.g., triangular or normal distributions based on expert judgment and historical variability). The specific ranges and distributions were derived from a range of possible real-world outcomes (e.g., slightly lower forecast accuracy than optimal, or less efficient optimization in unforeseen scenarios).
Figure 4 presents the resulting probability distributions for NPV and BCR, providing a probabilistic view of the project’s financial outcomes under uncertainty.
The Monte Carlo simulation indicates a mean NPV of approximately €365 million and a mean BCR of 1.41. The probability of achieving an NPV greater than zero is estimated to be over , while the probability of a BCR greater than 1.0 is also over . The slight difference between the Monte Carlo mean NPV/BCR and the deterministic base case values can arise from asymmetric probability distributions of input variables, the presence of significant non-linear relationships between input variables and the output metrics, or correlations between input variables, which are inherently captured in a probabilistic approach. These results strongly reinforce the financial robustness of the ODP investment under uncertainty, demonstrating high confidence in its viability even when considering potential variations in adoptions and discount rates, and other key parameters.
Qualitative risk assessment identified additional factors: regulatory and policy risks (e.g., changes in market design or data sharing regulations), cybersecurity and data privacy risks (given the sensitive nature of energy and mobility data), technology adoption and integration risks (e.g., slower than expected stakeholder uptake, challenges in integrating with legacy systems), and stakeholder acceptance. Mitigation strategies include proactive engagement with policymakers to shape supportive regulatory frameworks, robust cybersecurity measures adhering to international standards like ISO 27001 and GDPR compliance, a phased deployment approach to manage technical complexities, and a user-centric design philosophy to enhance stakeholder acceptance.
The sensitivity patterns observed in the aggregated analysis (Section 4) generally hold true at the country level. This indicates that while the absolute values of benefits and costs vary, the project’s overall viability and robustness to uncertainties remain consistent across different national contexts. Monte Carlo simulations performed independently for each country consistently indicate a probability of achieving a BCR 1 exceeding in all three national cases. This further reinforces the economic rationale for the ODP’s deployment at both regional and national scales, providing a high degree of confidence in its financial success despite inherent market and technological uncertainties.
5 Discussion
The comprehensive CBA conducted for the AI-driven ODP clearly indicates that the proposed platform offers substantial net benefits for the integrated management of electric mobility, renewable energy, and grid operations. The aggregated positive net present value (NPV) of €356.7 million and a benefit-cost ratio (BCR) of 1.41 provide a strong economic rationale, signifying that every Euro invested generates approximately €1.41 in benefits over the 10-year analysis period.
The main insight from this study is the substantial economic and operational value unlocked specifically by the ODP’s integrated AI engine. The largest share of the quantified economic benefits stems directly from AI-enabled functionalities, a linkage systematically established through specific ODP performance factors integrated into the benefit monetization (as detailed in Section 4 and A.1). For instance, the revenue from optimized energy trading and ancillary services (ROETAS) and cost savings from demand response and peak load reduction (CSDR-PLR) combined account for over 25.8% of the total benefits, directly attributable to the superior performance of the ODP’s AI forecasting and optimization models.
The case studies across Austria, Hungary, and Slovenia demonstrate the ODP’s adaptability and value proposition across varied national energy and transport landscapes. The substantial reliability benefits (€348 million in GSMS, alongside €171 million in CSDR-PLR as an economic benefit) underscore the ODP’s potential as a highly effective non-wires alternative, achieved through improved system visibility, AI-driven proactive control, and enhanced coordination across sectors. From a societal and policy perspective, the significant environmental benefits (€127 million from CO2 and air pollutant reductions) are critical, highlighting the ODP’s role in accelerating decarbonization. The robustness of AI models to real-world conditions such as data drift, concept drift, and unexpected events is planned to be managed through continuous performance monitoring, automated retraining pipelines, and robust data validation, which are all factored into the recurring OPEX (A.3), ensuring sustained long-term benefits.
While the comprehensive CBA provides a strong economic rationale, several limitations warrant acknowledgment. Firstly, while rigorous backtesting and validation were performed, and simulations incorporate realistic assumptions, there remains a potential for variance in real-world AI model performance versus simulated benchmarks, particularly under unforeseen extreme conditions. Future pilot projects are essential for further empirical validation. Secondly, the optimal functioning of the ODP’s AI engine relies on sustained, high-quality, and comprehensive data availability from a multitude of disparate sources. These include real-time EV and ET fleet telemetry, smart charging infrastructure, renewable energy forecasts, market data from EPEX Spot and ENTSO-E, weather models, and public datasets from Eurostat and the EEA.
Thirdly, certain intangible benefits, such as enhanced user convenience, improved public health outcomes from cleaner air that are difficult to fully monetize, or the empowerment of consumers in energy management, were not explicitly quantified in monetary terms, potentially leading to an underestimation of the true societal value. Additionally, the study makes assumptions regarding rational user behavior in response to ODP incentives and information, which may vary in practice. The dynamic nature of the regulatory landscape influencing energy markets, data sharing, and cross-border energy flows introduces inherent uncertainties. Finally, while the three selected EU member states represent a diverse range, there are inherent challenges in quantitatively generalizing the precise findings to contexts vastly different from those studied, although the qualitative benefits of integrated, AI-driven energy and transport management are broadly applicable globally.
6 Conclusion
This paper has presented a comprehensive cost-benefit analysis of an AI-driven operational digital platform developed for the integrated management of electric vehicles, electric trucks, renewable energy sources, and the power grid. The CBA framework was designed in alignment with European Commission guidelines for energy infrastructure projects and tailored to explicitly capture AI-specific value streams. These include improvements in forecasting accuracy, operational optimization, grid reliability, and emissions reduction, each monetized using transparent and reproducible methodologies. Importantly, the framework incorporates both capital and operational expenditures, particularly those linked to AI model development, deployment, and maintenance, alongside benefits categorized as economic, reliability, and environmental.
The analysis demonstrates that the proposed AI-driven ODP yields substantial net economic and societal value. With a projected net present value of approximately €356.7 million and a benefit-cost ratio of 1.41 over the 2026–2035 operational period, the investment case is strong. These results are driven by the ODP’s multi-layered architecture, where the AI core, featuring LSTM-based forecasting models, MILP-based optimization algorithms, and electric vehicle routing solvers, serves as a critical value enabler. These technologies jointly support enhanced energy and operational efficiency, predictive congestion mitigation, optimized demand response, and battery degradation reduction through health-aware charging strategies. The robustness of these benefits was confirmed through extensive stress-testing, including sensitivity analyses and Monte Carlo simulations.
To validate cross-border applicability, the CBA was applied to Austria, Hungary, and Slovenia, each reflecting different energy and transport profiles. The findings illustrate how national characteristics influence the distribution of benefits: Austria achieves substantial fleet efficiency gains and extended battery life; Hungary benefits from improved RES integration and reduced curtailment; and Slovenia sees value in grid flexibility and optimized ET logistics. These variations underscore the scalability and adaptability of the ODP to diverse regulatory, technical, and market contexts.
Beyond financial metrics, the study offers strategic insights. For policymakers, the results affirm the role of advanced digital infrastructure in achieving EU-wide decarbonization goals under REPowerEU and Fit for 55. The ODP presents a viable pathway to accelerate transport electrification, enhance grid reliability, and reduce emissions through AI-based coordination. For industry stakeholders, including DSOs, TSOs, fleet operators, and aggregators, the analysis provides a quantitative basis for investing in interoperable, intelligent systems capable of delivering both operational gains and new service opportunities. At the same time, the study acknowledges key limitations. The optimal functioning of the AI engine depends on sustained access to high-quality, heterogeneous data sources, such as real-time telemetry from EVs and ETs, smart charging infrastructure, market platforms (e.g., EPEX Spot, ENTSO-E), and weather models. Moreover, generalizing the results to other countries requires careful consideration of differences in data infrastructure, policy environments, and electrification maturity. Additionally, intangible benefits, such as improved public health from air quality gains or increased user convenience, were not monetized, likely leading to a conservative estimate of the platform’s true societal value.
In conclusion, the integrated management of EVs, ETs, RES, and the grid via AI-driven digital platforms is shown to be not only technically feasible but also economically and environmentally beneficial. The proposed ODP offers a compelling blueprint for next-generation energy–mobility convergence, supporting the emergence of smart, flexible, and low-carbon energy ecosystems. By holistically orchestrating distributed assets through AI, the platform bridges the gap between high-level policy goals and on-the-ground operational needs.
Future work should focus on several key areas to further validate and expand the ODP’s capabilities and impact. This includes the initiation of pilot projects for empirical validation of the ODP’s AI models and overall system performance in real-world operational environments, providing concrete data on benefits and operational challenges. A deeper exploration of advanced V2X (Vehicle-to-Everything) services, including vehicle-to-home (V2H) and vehicle-to-grid (V2G) applications beyond ancillary services, could unlock further synergies. Detailed distribution grid impact modeling, specifically for local voltage stability and transformer loading under high EV penetration and dynamic charging, is crucial. Research into advanced cybersecurity measures tailored for AI-driven energy platforms, including AI-specific threat detection and mitigation, is paramount given the critical infrastructure nature. Policy and market design innovation is essential to incentivize broad stakeholder participation and ensure fair value distribution across the ecosystem. Finally, extending the ODP concept to integrate other flexible resources like smart heat pumps, stationary battery storage, and industrial flexible loads could further enhance system-wide optimization and accelerate decarbonization efforts.
Acknowledgements
This research was supported by the BEGONIA project, which has received funding from the European Commission under grant agreement No. 01133306.
Appendix A Detailed cost-benefit analysis methodology and data
This appendix provides further details on the methodologies and data underpinning the CBA presented in the main paper, ensuring transparency and reproducibility of the quantitative analysis.
A.1 Benefit monetization formulas and key parameters
This section details the primary formulas and key parameters used for monetizing benefits in Step 4: Monetize benefits of the CBA, as discussed in Section 4. All monetary values are expressed in real terms (Euro 2025), and future cash flows are discounted to their present value using a social discount rate of . Each benefit’s realization is linked to the capabilities of the AI engine described in Section 3.1. In the following subsections, we present the formulas for calculating each benefit, along with the specific country-related assumptions, technological assumptions, and projected data.
A.1.1 Reduction in operational downtime (ROD) cost
The benefit from reduced operational downtime is calculated as the monetized value of avoided unserved energy due to improved grid reliability and faster fault recovery facilitated by the ODP’s AI-driven predictive maintenance and enhanced situational awareness. The benefit from reduced operational downtime is modeled as:
where is the baseline operational downtime cost222A conservative 24.2 million €/year for the three countries is taken in our calculations., is the energy self-efficiency improvement fraction333For this benefit calculation we assume and , and is the ODP-enabled downtime cost savings, reflecting faster fault recovery and improved grid reliability due to AI-driven predictive maintenance and situational awareness (Section 3.1). The product captures the adjusted downtime cost under improved self-efficiency, and multiplying by converts this into the annual benefit from avoided unserved energy, consistent with analogous smart-grid reliability improvements [53].
A.1.2 Revenue from optimized energy trading and ancillary services (ROETAS)
This benefit quantifies increased revenues and cost savings from optimized market participation due to the ODP’s AI forecasting and optimization capabilities.
where stands for the energy arbitrage revenue at hour under an ODP-enabled system and is the corresponding revenue without ODP. The hourly arbitrage revenue itself is calculated as where is the ODP-enabled flexible capacity (MWh), is the AI-identified price spread (€/MWh), and is the round-trip efficiency. To explicitly consider the efficiency improvement on RES, ET, and EV due to ODP444For our calculation, the EV/ET efficiency improvement due to ODP is assumed to be 5%, of total annual energy, and €/MWh. Moreover, the RES capacity factor CF represents typical full-load hours for renewables; we use a conservative value of 1,200 h/year., one may replace with where denotes the net uplift fraction and [€/MWh] is the assumed average electricity price. The term represents the annual baseline renewable-energy output, defined as , where [MW] is the installed RES capacity and [h/yr] is its full-load hours. Similarly, is the annual baseline EV + ET energy in MWh, with [MW] the aggregated EV + ET charging capacity and h/yr. The set lists all ancillary-service products. For each and each hour , [€/MWh] and [€/MWh] denote the ODP’s revenue and cost for service , respectively, while [€/MWh] and [€/WWh] denote the corresponding baseline values under non-ODP operation.
A.1.3 Cost savings from demand response and peak load reduction (CSDR-PLR)
Avoided costs from reducing system peak demand are realized through ODP’s AI-driven demand response (DR) and peak-load-shifting capabilities. We model the annual benefit as
Where the first term is benefits from cost savings from demand response, and the second term is benefits from peakload revenue. [€/MWh] is the unit cost saving from demand response (CSDR) per shifted MWh; [MW] is the installed RES capacity and [h/yr] its full-load hours, hence is the annual RES-generation baseline in MWh. The factor is the fraction of RES output that can be shifted via DR555, €/MWh, and in our calculations.. Similarly, [MW] is the aggregated EV charging capacity, [MW] is the total ET charging capacity, the terms and are the annual EV and ET baseline energy (MWh) multiplied by the shiftable fractions and .
In the second term, [€/MWh] denotes the peak-load reduction (PLR) revenue per shifted MWh, and is the fraction of total annual energy that can be shifted away from peak hours. We also have:
where is the annual volume (MWh) shifted out of peak hours. Thus, the second term represents revenue from peak-load reduction at peak-period prices.
A.1.4 Avoided energy curtailment (AEC)
Monetized value of RES utilized rather than curtailed, enabled by ODP’s AI-driven RES forecasting and flexible load matching. This benefit is calculated using the following formula:
where is the fraction of RES curtailment avoided666=1% and 50 €/MWh in our calculation., [MW] is the installed RES capacity, and [€/MWh] is the unit cost associated with avoided curtailment. Thus, gives the annual baseline RES generation (MWh), multiplying by yields the avoided MWh, and multiplying by converts to euros.
A.1.5 Fleet energy and operational savings (FES)
Reduced EV/ET operational costs are achieved through ODP’s AI-optimized routing and charging, which extend battery lifespan and improve energy efficiency. The total annual benefit is calculated as follows:
where is the overall ODP-enabled savings fraction777We consider 1%, 800 €/year, 1920 €/year in our calculation., and are the total number of EVs and ETs, respectively, and are the effective efficiency factors capturing reduced energy consumption and extended battery life, and and are the annual per-vehicle energy and battery-degradation savings under ODP. The factor consolidates multiple AI-driven improvements: battery-health optimization (reducing degradation by via SoC and C-rate management [36, 37, 38]) and routing efficiency gains (3–5% savings from E-VRPTW optimization). Thus, reflects a conservative 1% net reduction in total fleet operating cost when multiplying the baseline EV and ET savings terms.
A.1.6 Grid stability and management cost savings (GSMS)
Reduced DSO/TSO costs for grid stability arise from several ODP-enabled improvements (Section 3.1). We calculate the annual benefit as
where is the ODP-enabled forecasting improvement fraction (informed by smart-grid pilot results [53, 54])888We assume , , , , , in our calculations., is the total number of EVs, is the total number of ETs, [MWh/yr] and [MWh/year] are the annual charging energy per EV and per ET, respectively, and [€/MWh] is the unit fuel-saving value. The product captures savings from reduced fuel use due to improved timing of EV/ET charging. The term represents avoided operational-downtime costs: is the baseline downtime cost, [MW] is installed RES capacity, and is the RES full-load hours, so that multiplying by estimates the portion of downtime cost deferral attributable to RES-driven stability improvements.
Finally, is the ODP-enabled grid-stability savings fraction. Multiplying by where [€/MWh] is the market-price reduction rate, yields additional benefits from reduced imbalance and congestion charges when EV/ET charging shifts under ODP’s real-time control.
A.1.7 CO2 emission reduction benefits
Monetized GHG emission reductions arise from ODP’s AI-driven increase in RES utilization and optimized EV/ET charging (Section 3.1). We model the annual benefit as
where is the ODP-enabled reduction fraction999we take 12%, 0.4 MWh/year, 0.8 MWh/year,0.025 tonne CO2/MWh, [€/tonne CO2]=50 €/tonne, in our calculation., is the total number of EVs and [MWh/year per vehicle] is the annual energy equivalent per EV after AI-optimized charging; similarly, is the total number of ETs and [MWh/year per vehicle] is the annual energy equivalent per ET. The factor [tonne CO2/MWh] is the emission factor, and [€/tonne CO2] is the CO2 price following EU guidance [50]. Thus, gives the combined annual MWh associated with RES, EV, and ET operation; multiplying by yields total tonnes of CO2 avoided, and converts to euros, with capturing the the incremental reduction enabled by ODP.
A.1.8 Reduction in air pollutant (RAP)
Monetized health and environmental benefits from reduced local air pollutants arise when ODP’s AI enables cleaner generation, demand response, and transport (Section 3.1). We model the annual benefit as
where [kg/km] is the EV-air pollutant emission factor101010We take 0.000005 kg/km, 20,000 km/year, 20%, 1%, 10%, 40,000 km/year, 12%, 0.025 kg/MWh in our calculations., [km/yr] is the annual distance per EV, and is the total number of EVs, so that gives the total annual EV-related pollutant reduction in kilograms. Here, is the fraction of RES capacity initially considered each year for the emitting pollutants calculation, is the fraction of RES output shifted through DR, and is the grid decarbonization factor; therefore represents the annual MWh of RES energy shifted and decarbonized. Similarly, [kg/km] is the ET-air pollutant emission factor, [km/year] is the annual distance per ET, is the total number of ETs, and is the ODP-enabled emission reduction fraction, so that gives the ET-related pollutant reduction in kilograms. The term [kg/MWh] is the average emission factor for SO2, NO2, and PM10, making the annual kilograms of pollutants avoided by cleaner RES generation. Finally, [€/kg] is the unit cost of local air pollutants based on EU guidelines [56].
A.2 CAPEX breakdown
This appendix details the capital expenditure (CAPEX) components for the AI-driven ODP. The figures in Table 9 represent the discounted present value (PV) of costs over a ten-year operational period (2026-2035), aggregated for deployment across Austria, Hungary, and Slovenia. The breakdown distinguishes between ongoing hardware investments for expanding sensor networks (which are incremental annually) and initial one-time investments for the core technology platform. The AI-specific CAPEX includes initial development of AI models and the dedicated computing infrastructure required for their training and deployment.
| Cost category | Total (PV) |
| Hardware (sensors and edge devices - incremental) | |
| Incremental sensors for new EVs | 280.00 |
| Incremental sensors for new ETs | 120.00 |
| Incremental sensors for new RES installations | 71.65 |
| Subtotal hardware (incremental) | 471.65 |
| Core technology platform (initial investments and upgrades) | |
| Adaptive environmental sensor arrays (initial deployment) | 20.00 |
| Weather forecasting sensor infrastructure | 1.20 |
| Connectivity infrastructure (5G/LPWAN) | 3.00 |
| Communication protocols | 3.00 |
| Message queue systems (Kafka) | 3.00 |
| Databases and data lakes (setup) | 3.00 |
| Data integration middleware (FIWARE Context Broker) | 4.50 |
| API management gateway | 4.50 |
| Market integration middleware (EPEX SPOT APIs) | 4.50 |
| AI models for grid optimization (development and initial training) | 15.00 |
| Demand response tools (integration) | 6.00 |
| Route optimization engines (development and integration) | 9.00 |
| Weather forecasting algorithms (development and integration) | 3.00 |
| End-user mobile application development | 0.90 |
| Charging navigation systems integration | 3.00 |
| Comprehensive analytics tools development | 3.00 |
| Administrative dashboard development | 3.00 |
| Software deployment and system integration | 3.00 |
| Subtotal core technology platform | 89.60 |
| Total CAPEX (PV) | 561.25 |
1.
Note: Incremental sensor CAPEX reflects the cumulative discounted cost of equipping newly adopted EVs, ETs, and RES units annually with ODP-compatible sensors and edge devices. Core Technology Platform costs represent foundational one-time investments for the ODP core infrastructure and software components, including dedicated AI development. These figures are based on industry benchmarks for similar digital platform deployments in the energy sector, scaled for the three-country scope. Contingencies are embedded within these estimates.
A.3 OPEX breakdown
This appendix details the OPEX components for the AI-driven ODP. The figures in Table 10 represent the discounted present value (PV) of recurring annual costs over a ten-year operational period (2026-2035), aggregated for deployment across Austria, Hungary, and Slovenia. The Annual average PV provides an annualized perspective of these costs, calculated from the total 10-year discounted OPEX. Costs specifically related to sustaining and evolving the AI engine are explicitly considered.
| Cost category | annual Avg. PV |
| Cloud computing (infrastructure, VMs, storage, networking) | 3.9 |
| AI inference and data processing (real-time predictions, model serving) | 3.9 |
| Software maintenance and updates (platform, third-party libraries) | 2.2 |
| AI model maintenance, retraining, and performance monitoring | 2.5 |
| Grid and sensor maintenance (upkeep of physical assets) | 2.8 |
| Cybersecurity and threat monitoring (platform-wide) | 3.6 |
| Customer support and data analytics (stakeholders and users) | 1.9 |
| TSO/DSO and energy market integration (data feeds, API costs) | 2.8 |
| Grid expansion and new EV/ET integrations (operational, excl. CAPEX) | 3.2 |
| Network maintenance (5G/LPWAN) | 3.0 |
| Market trading fees (transactions, commissions) | 2.4 |
| Miscellaneous costs (training, compliance, licenses) | 2.8 |
| Labor costs (technical staff, data scientists, operational support) | 3.0 |
| Total annual average OPEX (PV) | 31.6 |
| Total 10-Year OPEX (PV) | 315.9 |
1.
Note: The Annual Average PV is derived from the total 10-year discounted OPEX. The distribution across categories is proportionally guided by detailed first-year operational cost assessments, encompassing all recurring expenses for platform operation, maintenance, and support. AI-specific OPEX (e.g., Cloud computing for AI inference, AI model maintenance) accounts for approximately 20% of total OPEX, crucial for sustaining the AI engine’s performance. Contingencies are embedded within these estimates.
A.4 Per-country CBA summary and Monte Carlo simulation (MCS) results
This appendix provides a summary of the final CBA outcomes and Monte Carlo Simulation (MCS) results for the AI-driven ODP, disaggregated for each of the three case study countries: Austria (AT), Hungary (HU), and Slovenia (SI). The numerical values presented here are the actual calculated results based on country-specific assumptions and scaled costs, demonstrating the national-level impacts derived from the detailed financial and operational models.
| Metric | Austria (AT) | Hungary (HU) | Slovenia (SI) |
| Deterministic results (Base Case) | |||
| Total benefits (PV) | 550.8 | 434.7 | 248.4 |
| Total costs (PV) | 390.4 | 310.8 | 176.0 |
| NPV (PV) | 160.4 | 123.9 | 72.4 |
| BCR | 1.41 | 1.40 | 1.41 |
| Monte Carlo simulation (Mean values from 10,000 Trials) | |||
| Mean NPV (PV) | 168.2 | 127.5 | 74.3 |
| Mean BCR | 1.43 | 1.42 | 1.42 |
| Dominant benefit categories (% of national total benefits) | |||
| Economic (ROD, AEC, FES, ROETAS, CSDR-PLR) | |||
| Environmental (CO2, RAP) | |||
| Reliability (GSMS) | |||
1.
Note: These numerical values are derived from detailed country-specific financial and operational models, reflecting the specific growth rates, market prices, and grid characteristics outlined in Table 5 and other appendices. The slight variations in BCR across countries reflect the unique cost structures and benefit profiles of each national context. Dominant benefit categories highlight the primary drivers of value in each country, influenced by their respective energy and transport system characteristics.
The Monte Carlo simulation results for each country consistently demonstrate a very high probability of positive NPV and BCR greater than 1, reinforcing the robustness of the ODP investment across diverse national contexts within the EU. The distribution of benefits varies slightly, with countries like Austria showing higher absolute economic benefits due to a larger EV/ET stock and higher energy prices, while Hungary benefits significantly from avoided RES curtailment given its ambitious solar expansion. Slovenia, despite its smaller scale, shows strong relative benefits driven by optimized cross-border logistics and effective grid integration.
Appendix B Annual projections and underlying data for the ODP case studies
This appendix provides a detailed overview of the annual projections for key parameters, monetized benefits, and costs specifically for the AI-driven ODP case studies across Austria, Hungary, and Slovenia. These projections form the basis for the discounted present values presented in the main paper and are derived from the comprehensive modeling framework and assumptions outlined in Section 3.
B.1 Annual EV/ET Stock and RES capacity projections
The adoption rates of EVs and ETs, along with the growth of RES capacity, are fundamental drivers of the ODP’s benefits and costs. Table 12 and Table 13 present the annual projected stock of EVs and ETs, and the installed RES capacity for each of the three countries over the 10-year analysis period (2026-2035). These projections are consistent with national energy and climate plans (NECPs), established industry forecasts, and historical trends, as referenced in Table 5.
| Year | EV Stock (thousands) | ET stock (thousands) | ||||
| Austria | Hungary | Slovenia | Austria | Hungary | Slovenia | |
| 2026 | 343.8 | 100.8 | 50.0 | 4.4 | 0.4 | 0.3 |
| 2027 | 398.8 | 118.8 | 57.7 | 4.9 | 0.5 | 0.3 |
| 2028 | 462.6 | 140.0 | 66.6 | 5.5 | 0.5 | 0.4 |
| 2029 | 536.6 | 165.1 | 76.9 | 6.2 | 0.6 | 0.5 |
| 2030 | 622.5 | 194.7 | 88.8 | 7.0 | 0.7 | 0.5 |
| 2031 | 722.1 | 229.6 | 102.6 | 7.9 | 0.8 | 0.6 |
| 2032 | 837.7 | 270.6 | 118.5 | 8.9 | 0.9 | 0.7 |
| 2033 | 971.7 | 318.9 | 136.9 | 10.1 | 1.0 | 0.8 |
| 2034 | 1127.2 | 375.9 | 158.2 | 11.4 | 1.1 | 0.9 |
| 2035 | 1307.5 | 443.3 | 182.7 | 12.8 | 1.3 | 1.0 |
| Year | Austria (GW) | Hungary (GW) | Slovenia (GW) |
| 2026 | 10.5 | 8.0 | 1.5 |
| 2027 | 11.2 | 10.7 | 1.6 |
| 2028 | 11.8 | 13.5 | 1.8 |
| 2029 | 12.5 | 16.3 | 1.9 |
| 2030 | 13.2 | 19.1 | 2.1 |
| 2031 | 13.9 | 21.9 | 2.2 |
| 2032 | 14.6 | 24.7 | 2.4 |
| 2033 | 15.3 | 27.5 | 2.5 |
| 2034 | 16.0 | 30.3 | 2.7 |
| 2035 | 16.7 | 33.1 | 2.8 |
B.2 Annual monetized benefits of the ODP (aggregated)
Table 14 presents the annual monetized benefits of the ODP, aggregated across Austria, Hungary, and Slovenia, prior to discounting. These figures demonstrate the year-on-year growth of benefits as the ODP’s functionalities mature and as EV/ET adoption and RES integration increase.
| Year | ROD | ROETAS | CSDR-PLR | FES | AEC | GSMS | CO2 | RAP | Total annual benefits |
| 2026 | |||||||||
| 2027 | |||||||||
| 2028 | |||||||||
| 2029 | |||||||||
| 2030 | |||||||||
| 2031 | |||||||||
| 2032 | |||||||||
| 2033 | |||||||||
| 2034 | |||||||||
| 2035 | |||||||||
| Total |
B.3 Annual costs of the ODP (aggregated)
Table 15 details the annual CAPEX and OPEX for the ODP, aggregated across Austria, Hungary, and Slovenia, prior to discounting. These figures illustrate the investment profile over the 10-year period, with initial higher CAPEX followed by recurring operational costs.
| Year | Annual CAPEX (€ million) | Annual OPEX (€ million) | Total annual costs (€ million) |
| 2026 | |||
| 2027 | |||
| 2028 | |||
| 2029 | |||
| 2030 | |||
| 2031 | |||
| 2032 | |||
| 2033 | |||
| 2034 | |||
| 2035 | |||
| Total |
1.
Note: Annual CAPEX and OPEX are discounted sums over the 2026–2035 period, aggregated across Austria, Hungary, and Slovenia. The Total row shows the ten-year aggregate costs for each category. In addition, a one-time upfront fixed cost of €89.6 million is incurred in 2026 to cover connectivity (5G/LPWAN), communication protocols and message queues, databases and data lakes, data integration middleware, market integration middleware, AI models for grid optimization, demand response tools, route optimization engines, weather-forecasting algorithms, software deployment and integration.