### v2

# Cost-benefit analysis of an AI-driven operational digital platform for integrated electric mobility, renewable energy, and grid management

Abstract
Integrating electric mobility, such as electric vehicles (EVs), electric trucks (ETs), and renewable energy sources (RES) with the power grid is paramount for decarbonization, efficiency, and stability. Two gaps remain: the need for digital platforms that coordinate operations across sectors and borders, and the lack of a rigorous framework on their cost–benefit profile. This paper presents a comprehensive cost–benefit analysis (CBA) of an AI-driven operational digital platform (ODP) designed for holistic, cross-sectoral and cross-border optimization. The ODP targets improvements in energy efficiency, grid reliability, and environmental sustainability. We develop a seven-step CBA framework that links each benefit stream to the platform’s architecture and quantifies incremental gains relative to the existing system, while accounting for investment and operating costs. The framework is demonstrated with case studies for Austria, Hungary, and Slovenia. Results indicate a benefit-cost ratio (BCR) of about 1.41 and a net present value (NPV) exceeding €356 million over 2026–2035. Robustness is examined through extensive sensitivity analyses that vary discount rates, cost components, and adoption trajectories, as well as through Monte Carlo simulations that capture uncertainty in BCR, NPV, data availability, and AI accuracy. The findings support the viability of investing in digital platforms for cross-sectoral and cross-border integration, and highlight the role of ODPs in advancing decarbonization and efficiency in the mobility–power nexus.
keywords:
Operational digital platform , Cost-benefit analysis , AI-driven optimization , Renewable energy integration , Transport electrification1 Introduction
The European Union is accelerating the shift to sustainable energy and transport, with rising adoption of electric vehicles (EVs) and electric trucks (ETs) and a growing share of renewable energy sources (RES) in the grid [Motlagh et al., 2025, International Energy Agency (2021), IEA, Börjesson and Proost, 2025, Breed et al., 2021]. The EU targets EVs/ETs to reach 60% of new vehicle sales by 2030 [Mądziel and Campisi, 2023]. The effective integration of burgeoning EV/ET charging demands with intermittent RES generation and existing grid infrastructure presents substantial technical and economic challenges [Yuvaraj et al., 2024, Psarros and Papathanassiou, 2023]. These include managing the spatiotemporal mismatch between variable renewable supply and dynamic charging demand, enhancing grid flexibility to accommodate new, potentially high-impact loads, and overcoming the lack of coordinated platforms for holistic management across the energy and transport sectors. Traditional approaches often operate in silos, leading to suboptimal asset utilization, grid congestion, curtailment of renewable energy, and underutilization of emerging flexibility resources.
The challenge becomes more pronounced for heavy-duty vehicles. Studies have shown that battery-electric propulsion is technically viable but economically constrained for long-haul freight applications due to battery weight, range limitations, and gaps in charging infrastructure [Çabukoglu et al., 2018, Cunanan et al., 2021]. Additionally, analyses comparing battery trucks to electric road systems (ERS) highlight the complexity of infrastructure investment decisions, especially in cross-border settings [Börjesson and Proost, 2025, Deshpande et al., 2023]. For instance, Börjesson and Proost [2025] emphasizes that without coordination, smaller countries may opt out of ERS investments, even if the overall EU welfare would benefit from collaboration, underscoring the need for integrated digital solutions that can bridge national and sectoral divides.
Cost-benefit analysis (CBA) provides a systematic framework for evaluating the economic, financial, and societal viability of complex integrated solutions like the one proposed herein. In European projects, standardized CBA methodologies have been developed to assess smart grid and sustainable energy projects. Notably, the European Commission Joint Research Centre (EC JRC) [EU Commission JRC, 2012] outlines a multi-step CBA approach tailored for smart grid investments in developing countries. On the other hand, while Stromsather et al. [2018] details a similar framework applied to a smart grid pilot project in Isernia, Italy, focusing on distribution system automation. Both methodologies advocate a sequential analytical process, encompassing scope definition, benefit identification, monetization, and uncertainty assessment.
Applying these CBA frameworks to comprehensively evaluate smart grid solutions that simultaneously manage energy and transport systems, particularly those leveraging high RES penetration, requires careful and detailed consideration of unique benefit streams and cost structures. While existing frameworks provide a robust foundation for smart grid solutions, there is no CBA framework for AI-driven operational digital platforms (ODPs) that rigorously quantify the specific contributions of AI functionalities and cross-sectoral and cross-border synergies. Even recent evaluations of integrated e-mobility systems often remain confined to national boundaries or siloed sectors [de Saxe et al., 2023, Greaker et al., 2022]. Studies assessing the economic viability of V2G strategies and smart charging [Greaker et al., 2022], or real-time optimization for EVs [Feng and Figliozzi, 2013, Vázquez-Canteli and Nagy, 2019], often focus on partial benefits and neglect the compounded synergies across sectors and borders that a unified, AI-driven digital platform could unlock. There is currently no CBA framework that utilizes modern AI and digital platforms to achieve cross-sector and cross-border integration and scalable smart technologies.
In this work, we propose a CBA framework to assess an ODP that integrates EVs/ETs, RES, and the grid. While many platforms focus on specific functions such as EV smart charging [Schott et al., 2019] or RES forecasting [Ahmed et al., 2020], the ODP in this work takes a broader view. It coordinates charging-station availability, dynamic routing for EV/ET fleets with energy and battery-health constraints, demand response programs, and renewable forecasting within one intelligent ecosystem. Specifically, this integrated management, orchestrated by the ODP’s AI engine, is designed to boost grid flexibility by dynamically aligning EV/ET charging patterns with RES availability. This, in turn, enhances overall system reliability through predictive load balancing and congestion management, yielding substantial economic benefits via optimized energy procurement, trading, and operational cost reductions, as well as potentially extending EV/ET battery lifespans. To quantify these outcomes, we apply a seven-step CBA framework, thoroughly informed by EU Commission JRC [2012] and Stromsather et al. [2018]. Although we follow EU guidelines for developing our CBA framework, our contribution extends beyond application to the development of a stylized, holistic methodology that explicitly considers the ODP core architecture. In this regard, there is no prior CBA framework that can be applied to operational digital platforms, given the significant investments required for integrating digital platforms and AI models. Additionally, we tested the framework with a practical case study that considered the interests of stakeholders qualitatively and the financial and economic feasibility quantitatively. Our application of this framework emphasizes the careful quantification of benefits directly attributable to the ODP’s capabilities and the synergistic effects of cross-sectoral integration, ensuring that AI-specific development, operational costs, and performance variabilities are explicitly addressed. Regarding the mobility dimension of the ODP, it is important to note that the proposed CBA framework is not an operational optimization of EV/ET logistics. Rather, it quantifies the incremental benefits of the AI-driven ODP by parameterizing EV/ET characteristics together with RES and grid interactions. Key factors include higher energy use per kilometer, reduced charging flexibility due to duty schedules. These parameters are incorporated into the benefit and cost channels (e.g., curtailment reduction, peak demand management, battery degradation savings), allowing EV/ET-specific constraints to influence system-level outcomes without requiring a full routing or scheduling model. This detailed application ensures alignment with EU practices while providing an investment decision-making CBA framework for AI-driven, multi-sector, and cross-border platform. Benefits are anchored to the ODP-enabled system compared with a system without ODP, using it as a baseline. Then, costs and benefits are quantified through platform-to-KPI linkages (e.g., curtailment reduction, peak shifting, imbalance reduction, congestion reduction, and reliability improvements). Environmental benefits valuation uses country-specific, time-resolved factors rather than static averages, and uncertainty analysis explicitly tests AI/data performance and adoption trajectories. This integrated, cross-sector, and cross-border framework offers a methodological advance beyond standard CBAs, which focus on single infrastructures, and yields a novel model for evaluating digital platforms that coordinate mobility, energy, and grid systems. We illustrate the CBA framework using a deployment of the ODP in Austria, Hungary, and Slovenia. Incorporating country-specific, technology, and economic assumptions, the case study reflects the cross-sector and cross-border nature of the ODP as it integrates diverse EV, ET, RES, and grid conditions. Therefore, the main contributions of the paper are threefold:
i)
Proposing a seven-step CBA framework to evaluate an AI-driven cross-sector, and cross-border ODP, explicitly identifying and monetizing AI-specific cost drivers and quantifying AI-driven benefit improvements (such as short-term RES forecasting and EV/ET charging optimization) as they propagate through the integrated transport–energy system.
ii)
Demonstrating a multi-layered, AI-driven ODP architecture by enumerating its core AI modules for RES forecasting, battery-health–aware EV charging, E-VRPTW routing for electric trucks, and co-optimization of grid ancillary services, and showing how verifiable performance metrics. This helps quantify economic, reliability, and environmental benefits, and the interconnectedness of each layer in its architectural core.
iii)
Applying the proposed framework to real-world case studies in Austria, Hungary, and Slovenia by utilizing country-specific input data and projections (CAPEX, OPEX, RES capacity factors, EV/ET fleet sizes and their projections, and wholesale price time series) and computing present-value streams for each benefit alongside CAPEX/OPEX. This gives actionable policy and managerial insights to inform Fit for 55 and REPowerEU objectives.
The remainder of the paper is structured as follows: Section 2 reviews related work. Section 3 presents the AI-driven ODP architecture and the proposed CBA framework. Section 4 presents the CBA steps with a case study and application. Section 5 discusses the main results of the study and some of the limitations. Finally, Section 6 offers conclusions and future directions.
2 Related work
The pursuit of sustainable energy transitions has spurred significant research into the application of AI for optimizing the complex interplay between electric mobility, RES, and power grid operations [Huang et al., 2025, Motlagh et al., 2025]. On the one hand, increasing reliance on digital platforms and smart charging programs can increase the risks of cyberattacks and the impact of natural disaster on power systems [Diaba et al., 2024]. The integration of EVs/ETs into the power system, which is overwhelmingly dominated by RES, has been the focus of several recent comprehensive reviews. Ref. Peng et al. [2012] analyzes the economic dispatch and risk management challenges posed by large-scale plug-in hybrid EV/ET integration into power systems, emphasizing the need for coordinated scheduling strategies and robust optimization under uncertainty. Similarly, Shariff et al. [2022] provides a state-of-the-art review on the impact of fast EV charging on the utility sector, highlighting critical concerns such as load spikes, grid congestion, and the lack of regulatory readiness for widespread EV adoption. Flexibility needs in power systems with high levels of EV penetration and RES are reviewed by Kaushik et al. [2022], who underscores the importance of adaptable system architecture to ensure reliability under dynamic conditions. On the other hand, recent literature highlights the transformative potential of AI in these domains, providing a robust academic foundation for the development and evaluation of integrated solutions, such as the AI-driven ODP proposed in this paper [Zhang et al., 2025, Child and Breyer, 2017, Rahman et al., 2022, Kumar et al., 2019]. These studies affirm that AI enhances the management of EVs, facilitates deeper RES integration, and enables more efficient and resilient grid operations, leading to considerable economic and environmental benefits. While the consensus points towards the positive impact of AI, the literature also acknowledges practical challenges such as data quality, system interoperability, cybersecurity, and the scalability of AI solutions, which are critical considerations addressed in the ODP’s design.
A foundational overview of AI’s role in the energy transition is provided by Ref. Wang et al. [2025]. The review highlights AI’s contributions to RES deployment and energy-efficiency improvements. It also covers enhanced grid stability via smart-grid integration and the management of demand-side dynamics, including vehicle-to-grid (V2G) systems. Their findings directly support the ODP’s core objective of orchestrating EVs, ETs, RES, and grid operations through sophisticated AI-driven forecasting and optimization. Complementing this, the survey by Ref. Al Nasim et al. [2024] on AI in the renewable energy domain identifies key AI-based strategies, such as neural networks and advanced optimization algorithms, that enhance the integration of diverse renewable energy sources, such as solar and wind. This broad perspective informs the selection and design of the ODP’s AI components, including its Long Short-Term Memory (LSTM) models for forecasting and mixed-integer linear programming (MILP) algorithms for optimization, with the goal of achieving robust and scalable solutions.
Focusing specifically on electric mobility, the systematic reviews by Arévalo et al. [2024], Rahman et al. [2022], Arias-Londoño et al. [2020] on the integration of AI into EV energy management systems reveal significant advancements. Ref. Arévalo et al. [2024] analyzes numerous studies that highlight progress in energy optimization, intelligent route planning, charging demand forecasting, and predictive maintenance for EVs, primarily leveraging machine learning, deep learning, and genetic algorithms. On the distribution network level, Ref. Rahman et al. [2022] conducts an in-depth impact analysis of integrating projected EV charging loads into low-voltage systems, revealing significant concerns about voltage stability and transformer loading. Earlier work by Ref. Arias-Londoño et al. [2020] also provides a chronological review of EV/ET interactions with distribution systems, laying out both technical and regulatory gaps. The anecdotal analyses of AI techniques, by enhancing EV performance and battery health, directly underpin the ODP’s smart charging functionalities and logistics optimization modules, which notably incorporate considerations for battery degradation.
The challenge of integrating large-scale RES into existing energy systems is another critical focus area where AI demonstrates considerable promise. Ref. Liu et al. [2022] investigates AI’s pivotal role in managing RES within multi-energy systems, particularly for achieving carbon neutrality. They emphasize AI’s capacity for intelligent operational control, achieved through precise forecasting of RES generation and effective demand-response strategies. This aligns closely with the ODP’s objective of synchronizing EV/ET charging demand with intermittent RES availability, thereby minimizing curtailment and maximizing the utilization of clean energy. The practical applications and identified limitations discussed by Liu et al. [2022], such as the need for sophisticated and adaptive AI methodologies, provide valuable insights that have shaped the design of the ODP’s AI engine.
Parallel to advancements in RES integration and e-mobility, AI is recognized as a key enabler for the fundamental transformation of electrical grids into smarter, more resilient, and sustainable energy infrastructures. Ref. Rajaperumal and Columbus [2025] explores this transformative impact, tracing the evolution from centralized to increasingly decentralized grid structures. Their work highlights AI’s crucial contributions to real-time data analytics, predictive maintenance of grid assets, and the optimization of demand-response programs. These AI-driven capabilities significantly enhance overall grid efficiency, operational flexibility, and resilience. This directly corresponds to the ODP’s functions of improving grid stability and reliability through proactive, intelligent control strategies. The frameworks discussed for managing decentralized networks and ensuring interoperability inform the design of the ODP’s middleware layer and its approach to facilitating cross-border energy coordination. Such grid transformation efforts naturally lead to questions of economic justification.
Consequently, the economic and environmental viability of AI-driven smart energy systems becomes a crucial aspect for their practical implementation. Ref. Gudlaugsson et al. [2023] addresses this issue in its analysis of the InteGRIDy project, by conducting a detailed cost and environmental benefit assessment of integrating renewable energy and smart technologies. Their study quantifies tangible economic savings and significant environmental benefits, such as reductions in CO2 emissions. This work provides a pertinent methodological precedent and empirical support for the comprehensive seven-step CBA framework employed in our evaluation of the ODP.
This body of research highlights the significant potential of AI to revolutionize energy and transport systems. Despite individual studies demonstrate clear benefits in specific domains such as EV management, RES forecasting, or grid optimization, a clear research gap remains for a holistic digital platforms that integrate mobility and energy sectors. This study, therefore, proposes a novel CBA framework employing a comprehensive AI-driven ODP architecture for a cross-sector and cross-border energy-transport efficiency. The main objective is quantifying the incremental costs and benefits derived from the integrated AI functionalities, an aspect often treated qualitatively or with a narrower scope in prior work. The proposed CBA framework not only integrates EVs/ETs, RES, and grid, but also conducts a detailed CBA to quantify the potential benefits and expenditures.
3 AI-driven ODP and the CBA framework
This section presents the AI-driven ODP system architecture, an integrated platform for EV, ET, RES, and grid coordination, followed by the proposed CBA framework used to assess the financial and economic feasibility of the investment. Section 3.1 describes the multi-layered ODP architecture and AI engine, while Section 3.2 introduces the structured seven-step CBA framework.
3.1 ODP system architecture and AI engine
The AI-driven ODP is conceptualized as a multi-layered, interoperable platform designed to integrate a diverse array of technologies and stakeholders across the energy and transport domains. Its architecture, detailed in Table LABEL:tab_architecture, is structured into five principal layers:- Perception, Data, Middleware, Business, and Application. This layered design promotes modularity, scalability, and maintainability. The robustness of AI components is ensured through continuous monitoring, periodic retraining pipelines to counter data/concept drift, and leveraging scalable cloud infrastructure for computational needs, ensuring high availability and performance even with increasing data volumes and model complexity.
| Layer | Requirement | Type of technology / AI component | Main required technologies / AI model specification |
[…]
1.
Note: The ODP architecture is based on the EU-funded BEGONIA project; all layers are described in detail in Deliverable D3.1 Mombartz-Brændgaard et al. [2024].
Note that the scope here is to assess the economic and financial viability of the proposed AI-driven ODP via a transparent CBA framework. This means we do not empirically validate or optimize the AI modules or the core architecture presented in Table LABEL:tab_architecture. Accordingly, any performance figures (e.g., forecasting accuracy or coordination gains) are treated based on the literature values of the platform when deployed in the implementation phase, rather than as case-study results. The requirements and benefits of each architectural layer should be read together with the technological diagram in Figure 1 to understand the platform-level interconnectedness. The CBA then evaluates the economic and financial benefits and expenditures required to implement this ODP using the case study. Below are the details of each layer:
The Perception layer is responsible for real-time data acquisition from a multitude of distributed assets, including EVs, ETs, RES installations (solar PV, wind turbines), charging infrastructure, and environmental sensors. This layer employs various sensors (BMS data, GPS, smart meters, pyranometers and anemometers) and communication technologies (5G, LTE-M, LoRaWAN) to ensure robust and timely data transmission. Data is collected at granular resolutions (e.g., 1-minute intervals for critical parameters).
The Data layer manages the ingestion, processing, storage, and secure handling of these voluminous and heterogeneous data streams. It leverages communication protocols (MQTT, OPC-UA), message queuing systems (Apache Kafka), and a hybrid data storage architecture comprising time-series databases (InfluxDB, TimescaleDB) for operational data and data lakes (e.g., Azure Data Lake) for large-scale analytics and AI model training. Data quality validation, anomaly detection, and missing data imputation are integrated processes within this layer to ensure robust inputs for the AI engine.
The Middleware layer provides crucial interoperability services, enabling seamless integration with existing utility systems (SCADA/DMS/EMS), energy market platforms (e.g., EPEX SPOT, Nord Pool APIs), and emerging EU data spaces (e.g., Gaia-X, IDS connectors). It utilizes enterprise integration patterns, API management gateways, and connectors designed for specific data exchange standards (e.g., ENTSO-E data models, Open Charging Protocols). This layer ensures secure, standardized, and real-time data flow necessary for cross-sectoral coordination.
The Business layer constitutes the core of the AI engine in the ODP, housing advanced algorithms and models that drive intelligent decision-making. Key AI components include:
Forecasting models: Long Short-Term Memory (LSTM) based Recurrent Neural Networks (RNNs), enhanced with multi-head attention mechanisms, are employed for forecasting EV/ET charging demand, RES generation, and grid load at various temporal and spatial resolutions (e.g., 15-minute intervals up to 72 hours ahead). These models capture complex temporal dependencies and non-linear relationships inherent in energy data, using inputs, such as historical load/generation, high-resolution weather forecasts (temperature, irradiance, wind speed, cloud cover from numerical weather prediction models, and local sensors), calendar effects, and real-time market signals. Model training is performed on extensive historical datasets, demonstrating a significant uplift in predictive power essential for subsequent optimization steps. Energy price forecasting models with machine learning for residual analysis, are alsoincluded in the core ODP architecture to predict day-ahead and intra-day electricity prices.
EV/ET smart charging and V2X dispatch: for this service, mixed-integer linear programming (MILP) or mixed-integer quadratic programming (MIQP) formulations optimize charging/discharging schedules for individual vehicles, and aggregated fleets can be augmented. Within the ODP, smart charging and V2X coordination are treated as benefit channels. The platform schedules charging and limited vehicle-to-grid export to align with RES availability, tariff signals, and grid limits, while respecting operational requirements (e.g., departure times and minimum state-of-charge). Any battery-health safeguards (e.g., avoiding extreme SoC or excessive C-rates) are handled as practical constraints rather than as a separate modeling contribution. The CBA captures the incremental value of this function through measurable KPI deltas, such as reduced peak demand, avoided curtailment, lower imbalance costs, and modest maintenance savings, relative to a clearly defined without-ODP baseline. In relation to the ODP AI-engine, the V2G/V2B services are managed by reinforcement learning (RL) agents (e.g., Deep Q-Networks or Proximal Policy Optimization). These agents are trained on high-fidelity simulations of real-time energy markets (reflecting bidding rules, settlement periods for frequency regulation services) and grid conditions. The RL reward function incentivizes revenue from ancillary services while safeguarding battery health and user mobility needs. Simulations incorporate historical market data, realistic network constraints, and stochastic elements to approximate real-world uncertainty.
ET route and logistics optimization: Sophisticated vehicle routing problem (VRP) algorithms, specifically tailored for electric fleets (E-VRPTW – Electric VRP with time windows and charging stops), are implemented. These algorithms optimize routes to minimize total operational cost (combining energy consumption, time, and potential battery degradation costs) while considering real-time traffic conditions, terrain (for energy consumption estimation), payload, dynamic weather conditions, charging station availability, charging duration, and potential queuing times at charging points. The integration with battery degradation models ensures that routing decisions do not inadvertently lead to accelerated battery wear.
Demand response (DR) management: AI-driven tools automate the identification and aggregation of flexible EV/ET loads, predict their response potential based on historical patterns and current vehicle states, and manage their participation in DR events or ancillary service markets. This adheres to standards like OpenADR 2.0b. The system utilizes AI-driven baseline calculation methods to accurately measure the effectiveness of DR events and optimize dispatch signals for individual vehicles or aggregated charging stations.
These AI models are developed using frameworks such as TensorFlow, Keras, and PyTorch, integrating with optimization solvers, such as CPLEX or Gurobi via Python APIs. Model training utilizes high-quality historical and real-time data from the Data layer. Crucially, the ODP incorporates continuous operational monitoring, performance validation metrics (e.g., forecast accuracy, optimization solution quality, solver runtime), and scheduled retraining pipelines to proactively address potential data drift, concept drift, or model degradation over time, ensuring sustained AI model efficacy and robust decision-making. Cloud-native scalable infrastructure supports the computational demands of large-scale AI model training and real-time inference.
Finally, the Application layer delivers user-facing interfaces and services. This includes intuitive mobile and web applications for EV/ET drivers and fleet managers (e.g., charge planning, station booking, real-time route guidance with charging recommendations), as well as comprehensive analytics dashboards and administrative consoles for system operators, charging point operators (CPOs), and energy providers. These dashboards provide transparent insights into system performance, AI model outputs, and optimization outcomes. This layered, AI-centric architecture allows the ODP to ingest vast quantities of real-time data, transform it into actionable intelligence through its AI engine, and orchestrate coordinated actions across the transport and energy sectors. The platform’s design inherently facilitates interoperability, dynamically aligning EV/ET charging with RES availability and grid conditions, while integrating market signals to optimize economic, environmental, and battery health outcomes, thereby creating an intelligent nexus for sustainable energy and mobility.
Figure 1 illustrates the interconnection of the five principal layers of the ODP. The perception layer collects real-time data from EV, ET, and RES sensors, as well as environmental and weather monitoring devices, which are transmitted via communication protocols and connectivity solutions. The middleware layer manages this data through message queues, ETL pipelines, and integration middleware, enabling secure exchange with external data spaces and market platforms. Processed data are then stored in data lakes and databases, forming the foundation for the service layer, where AI models for grid load, weather forecasting, demand response, and route optimization operate. The outputs feed into user-facing services such as mobile applications, dashboards, and analytics tools, which constitute the business layer by supporting operational decisions, market integration, and stakeholder engagement. This layered and interconnected design ensures scalability, interoperability, and coordinated optimization across EVs, ETs, RES, and the grid.
While the ODP–CBA relies on country-specific inputs (RES generation/forecasts, EV/ET fleets and charging schedules, grid constraints and states, cross-border flows), these data are increasingly standard in Europe due to smart-grid deployments, national data hubs, and harmonized interfaces. The framework is designed to plug into these interoperable data rails, making large-scale deployment practical where electrification and data-sharing initiatives are underway.
3.2 Cost-benefit analysis framework
To evaluate the proposed AI-driven ODP, we employ a structured seven-step CBA framework, inspired by the European Commission smart grid guidelines [EU Commission JRC, 2012] and lessons from pilot studies such as Isernia [Stromsather et al., 2018]. The proposed framework systematically captures AI-related costs (platform development, sensors and IoT devices, model training and maintenance) and the benefits directly attributable to AI-driven performance (e.g., improved forecast accuracy, and optimization savings). The framework begins with a precise definition of the project, its objectives, engineering features, geographic scope, stakeholders, and expected socio-economic impact, and delineation of its functional scope. Next, it maps the enabling technologies (especially AI modules) to the key functions they perform, and links those functions to the spectrum of benefits they generate. Subsequently, each category of benefit is quantified and monetized, and all associated costs (capital and operating) are tallied over the project’s life cycle. Finally, total benefits and costs are compared, and a rigorous sensitivity analysis, including scenario variations and Monte Carlo simulation, is conducted to assess how results vary under key uncertainties. Table 2 summarizes these steps.
| Step | Procedure |
| 1 | Define the project (scope, objectives, engineering features, AI functionalities, and their specific role). |
| 2 | Map technologies (including AI modules and their specific functionalities) to platform functions. |
| 3 | Map platform functions to economic, reliability, and environmental benefits. |
| 4 | Monetize and quantify benefits (linking ODP performance to benefit magnitude, with transparent assumptions). |
| 5 | Quantify costs (CAPEX and OPEX, including detailed AI system development, training, and maintenance costs). |
| 6 | Compare costs and benefits (NPV, BCR). |
| 7 | Conduct sensitivity analysis and risk assessment (including discount rate, technology adoption variability parameters, and Monte Carlo simulation). |
With this framework, Step 1 establishes the project’s scope, objectives, geographic coverage, stakeholders, and counterfactual baseline. Step 2 links each ODP technology layer and AI component to its operational function (data acquisition, integration, forecasting & optimization, and decision support). Step 3 identifies how these functions generate economic, reliability, and environmental benefits. In Step 4, each benefit is converted into monetary terms over the project horizon using transparent formulas and country-specific data. Step 5 aggregates all capital and operating expenditures necessary for deployment and operation, including AI model development, sensors, cloud services, and personnel. Step 6 computes net present value (NPV) and benefit–cost ratio (BCR) by discounting all cash flows, reflecting societal time preferences. Finally, Step 7 tests the robustness of the results through scenario variations and Monte Carlo simulation, and identifies key risks and corresponding mitigation strategies.
4 Cost-benefit analysis steps with case study
We now apply the seven-step CBA framework to the AI-driven ODP project, evaluating its deployment across Austria (AT), Hungary (HU), and Slovenia (SI). For the case study, i) the choice of the three countries is based on stakeholder analysis, the cross-sectoral and cross-border nature of the digital platforms. ii) The chosen countries represent different RES projection capacities, EV/ET penetrations, and input parameter prices. For data and parameter values, three assumptions are considered: iii) country-specific assumptions, technology-specific assumptions, and economic assumptions. All input assumptions, input parameters, and sources are provided in Tables (5, 10, 10, and 13). For the economic appraisal, all future cash flows are expressed in constant 2025 Euros and discounted to present value using an economic discount rate of 4%. [European Commission Directorate-General for Regional and Urban Policy, 2021]. The project’s primary objective is to use real-time data and advanced AI models to optimize system operations. In doing so, it will enhance grid flexibility, lower energy procurement and operating costs, cut greenhouse-gas and local-air-pollutant emissions, improve power-system reliability, and enable seamless interoperability between transport electrification and grid operations.
Step 1: Define the project
The project under consideration is the deployment and operation of the AI-driven ODP for the integrated management of EVs, ETs, RES, and the electric grid across three countries. The ODP’s primary objective is to use real-time data and advanced AI models to optimize system operations. In doing so, it will enhance grid flexibility, lower energy procurement and operating costs, cut greenhouse-gas and local-air-pollutant emissions, improve power-system reliability, and enable seamless interoperability between transport electrification and grid operations. The implementation scale is national/regional within these countries, which have been chosen for their diverse EV/ET adoption rates, renewable energy capacities, and grid characteristics, thereby providing a representative range of EU contexts for analysis.
Key engineering features of the ODP critical to the project include: i) AI-powered predictive algorithms for demand forecasting (EV/ET charging, overall grid load), RES generation forecasting, and energy price forecasting. ii) Optimization engines for smart charging scheduling, V2X services, electric fleet logistics (routing and charging), and demand response dispatch. These engines optimize multi-objective functions, including cost reduction, emissions reduction, and battery health preservation, thus enabling coordinated control strategies that balance the grid and maximize economic returns. iii) IoT infrastructure comprising a dense network of sensors and smart meters to provide granular, real-time monitoring of vehicles, charging stations, grid nodes, and RES installations. This high-resolution data collection is foundational for the AI engine’s situational awareness. iv) High-speed, reliable communication networks ensuring low-latency data transfer and control signal dissemination. This is critical for real-time AI inference and control, allowing the ODP to rapidly respond to changing conditions. v) Robust middleware integration with existing energy management systems (utility SCADA/DMS/EMS) and market platforms. By using standardized interfaces and data exchange protocols, the ODP can seamlessly interoperate with legacy systems and participate in energy markets, thus embedding the platform within the broader energy ecosystem.
As a stakeholder engagement, we conduct structured interviews and workshops with TSOs/DSOs, CPOs, fleet operators, and regulatory authorities to elicit requirements and validate governance, relevance, and technical aspects of the ODP. These assessments informed functional and data-sharing needs, adoption and usage assumptions, and implementation risks spanning governance, interoperability, and regulatory readiness.
Figure 2 summarizes a stakeholder-informed assessment across regulatory, technical, governance, economic/financial, and relevance dimensions for the three case-study countries, providing contextual evidence on organizational readiness and adoption feasibility that complements the quantitative CBA.
The benchmark for the CBA framework is the without-ODP system representing the existing system. The current system is largely uncoordinated in terms of EV/ET charging, has minimal V2G capabilities, and involves residual RES curtailment during surplus periods, as well as limited operator visibility/control due to legacy forecasting and tools. All reported benefits are then incremental relative to this benchmark. For benefit stream at time ,
where denotes the physical metric that is monetized and is annual benefits aggregate over and across countries. Assumptions regarding adoption, charging patterns, RES integration, and market conditions were informed by surveys and interviews with TSOs, DSOs, CPOs, fleet operators, and regulators in Austria, Hungary, and Slovenia.
Participation by EV drivers and fleet operators in smart charging and demand response is modeled explicitly as a scenario input. The CBA assumes rational, benefit-maximizing behavior within country-specific policies and grid codes (EV/ET, RES, network operation). Non-technical barriers (heterogeneous charging habits, DR fatigue, data-sharing and regulatory frictions) are not modeled as separate processes; instead, they are reflected through conservative/central/optimistic participation ranges and assessed qualitatively in the stakeholder-informed multi-criteria evaluation (regulatory, governance, relevance; Figure 2).
Step 2: Map technologies to functions
In this step, each core technology component of the ODP, particularly its AI modules as detailed in Section 3.1, is systematically mapped to the specific operational functions it enables. The ODP integrates a comprehensive suite of technologies, which together deliver several key functions for the system’s operation:
(A) Advanced data and information collection & processing: Real-time gathering, validation, and processing of granular data from EVs/ETs, RES installations, grid sensors, charging infrastructure status, and external sources (weather, market prices). This function is enabled by IoT sensors, connectivity solutions, and the ODP’s Data layer, ensuring that the AI engine has accurate and timely information about the system’s state at all times.
(B) Enhanced interoperability and system integration: Seamless data exchange and coordinated operation among diverse systems and stakeholders (vehicles, grid operators, CPOs, energy markets, end-users). This is facilitated by the Middleware layer and use of standardized communication protocols and interfaces. Enhanced interoperability ensures that decisions made by the AI core can be effectively communicated and executed across different subsystems in real time.
(C) AI-driven forecasting, optimization, and control: This core function is where the ODP’s AI engine directly drives value. It has two sub-functions:
i) Intelligent forecasting: The AI models (LSTMs, ML models, etc.,) provide highly accurate predictions for EV/ET charging demand, RES generation, grid load, and electricity prices. These forecasts improve upon traditional methods, reducing uncertainty and prediction error in the system’s operation.
ii) Smart load management and grid balancing: AI optimization algorithms (MILP, RL, E-VRPTW) enable dynamic smart charging/discharging of EVs/ETs, optimal dispatch of V2X services for grid support, dynamic EV/ET routing considering energy and battery health, and coordinated demand response actions. These control decisions continuously balance supply and demand, flatten load curves, alleviate grid constraints, and maximize the use of renewable energy.
(D) Decision support and automation: User-friendly interfaces (mobile apps, dashboards) provide actionable insights and recommendations to end-users (e.g., optimal charging times) and system operators (e.g., congestion alerts and recommended mitigation actions). Automated control systems execute optimized schedules and dispatch signals directly generated by the AI engine, ensuring real-time responses and reducing manual intervention.
Table 3 provides a high-level mapping of the ODP’s major technological assets and AI components to these functional categories A–D. For instance, IoT sensors and connectivity primarily support data collection (Function A). The integration middleware and communication protocols enable Function B. The AI forecasting models and optimization algorithms directly drive Function C (intelligent forecasting and control), while the user interfaces and automated control modules deliver Function D (decision support and action). This detailed mapping ensures that all critical functions necessary for the ODP’s success are supported by specific, well-defined technological capabilities, with the core AI engine acting as the central intelligence that orchestrates these functions.
| Technology (asset) / AI component | A | B | C | D |
| IoT sensors (EV, ET, RES, weather, grid) | ✓ | |||
| Edge gateways | ✓ | ✓ | ||
| 5G / LTE-M / LPWAN connectivity | ✓ | ✓ | ||
| Message queue systems (Kafka) | ✓ | ✓ | ||
| Communication protocols (MQTT, OCPP, OpenADR) | ✓ | ✓ | ||
| Microservice architecture | ✓ | ✓ | ||
| Data integration middleware (FIWARE) | ✓ | |||
| Data lake & time-series databases | ✓ | ✓ | ||
| AI forecasting models (LSTM, etc.) | ✓ | ✓ | ||
| AI optimization algorithms (MILP, RL, E-VRPTW) | ✓ | ✓ | ||
| Demand response management tools | ✓ | ✓ | ✓ | |
| External APIs (market, grid, weather) | ✓ | ✓ | ✓ | |
| User interfaces/dashboards (mobile apps, web) | ✓ | |||
| Automated control modules | ✓ | ✓ |
This systematic mapping confirms that the ODP’s technology stack, with its integrated AI engine, comprehensively addresses the functional requirements for data acquisition, system integration, intelligent analysis and control, and user interaction. This forms a robust foundation for optimizing energy and mobility services, creating synergistic benefits by enabling real-time, coordinated actions across the energy and transport domains that are not possible with siloed systems.
Step 3: Map functions to benefits
This step identifies the specific categories of benefits generated by each core ODP function. The anticipated benefits are categorized into three primary domains:
i) Economic benefits: Direct financial gains or cost savings, encompassing reduced operational expenditures for grid operators and end-users, new revenue streams from energy services, enhanced energy efficiency, and optimized asset utilization.
ii) Reliability benefits: Improvements in power system reliability and resilience, leading to reduced outage costs for consumers and businesses, enhanced service quality for end-users, and potentially deferred or optimized investments in grid infrastructure due to better load management.
iii) Environmental benefits: Reductions in greenhouse gas (GHG) emissions (primarily CO2) and local air pollutants (SOx, NOx, PM2.5), contributing to climate change mitigation, improved public health, and greater integration of clean renewable energy.
Each ODP function (A-D) contributes to one or more of these benefit categories. For example, advanced data collection and processing (A) provides the necessary real-time data for precise grid state estimation and load forecasting, enabling more efficient asset utilization and informed decision-making, which directly translates to economic efficiency and enhanced system reliability. Enhanced interoperability (B) enables dynamic coordination between different market participants and grid entities, enabling new revenue streams from flexible assets and reducing system imbalances. The core AI-driven forecasting, optimization, and control functions (C) are pivotal in generating significant economic benefits (e.g., reduced energy procurement costs through peak load shaving and optimized energy trading facilitated by highly accurate price predictions, and increased revenues from ancillary service provision). These AI functions also yield substantial reliability benefits (e.g., prevention of grid overloads, improved voltage/frequency stability through predictive control, and reduced RES curtailment by matching demand with variable generation). Furthermore, these AI functions directly contribute to environmental benefits by maximizing the use of RES, optimizing EV charging for lower emission periods, and reducing fossil fuel-based generation for balancing. Decision support and automation (D) translates AI-generated insights into tangible outcomes, yielding direct economic benefits for users (e.g., lower charging costs due to optimized schedules, time savings from efficient routing) and operators (e.g., automated response to grid events), and indirectly fostering environmental benefits by promoting behaviors aligned with RES availability and lower carbon intensity.
Table 4 qualitatively summarizes how the ODP’s key functions map to specific benefit mechanisms.
| Function | Economic | Reliability | Environmental |
| A. Advanced data & info. collection/processing | ✓ | ✓ | (i) |
| B. Enhanced interoperability/system integration | ✓ | ✓ | ✓ |
| C. AI forecasting, optimization & control | ✓ | ✓ | ✓ |
| D. Decision support & automation | ✓ | (i) | ✓ |
This mapping highlights that economic benefits are pervasively driven by all functional areas, with particularly strong and direct contributions from the AI engine (Function C) and effective user/system engagement through transparent decision support (Function D). Reliability benefits primarily stem from improved situational awareness (Function A), seamless integration enabling coordinated responses (Function B), and intelligent grid management capabilities (Function C) that anticipate and mitigate grid issues. Environmental benefits accrue mainly from functions that facilitate higher RES penetration and smarter, cleaner energy consumption patterns for EVs/ETs (Functions B, C, and D), minimizing reliance on fossil fuels.
Step 4: Monetize benefits
As described in Step 1, all benefits of the ODP-enhanced system are measured as the difference between the system with ODP and a counterfactual baseline (without ODP). This is done by defining a physical delta relative to the baseline and applying a valuation.
where is a monetary valuation of the delta-benefit, such as price.
The CBA adopts a 10‐year appraisal window (2026–2035) appropriate for a AI‐centric ODP whose enabling technologies undergo major upgrades within 5–10 years. This aligns with prior digital/smart‐grid CBAs Vilaplana et al. [2025], Ruutu et al. [2017] and avoids speculative long‐term assumptions typical of structural assets. A 15‐year sensitivity (2026–2040) using the same trajectories—with a mid‐life refresh—confirms the robustness of conclusions and yields higher NPV and BCR. Specific metrics, mathematical formulations, and assumptions are established for each benefit category. The details are presented in A.1, based upon standard valuation methodologies, relevant literature [EU Commission JRC, 2012, BloombergNEF, 2022], and country-specific data. For each relevant benefit, the improvement margin over the baseline is directly tied to the enhanced capabilities, either superior forecasting accuracy or more efficient operational optimization, provided by the specific AI components of the ODP. To demonstrate the ODP’s impact in varied national contexts, the proposed CBA framework is applied to Austria (AT), Hungary (HU), and Slovenia (SI). These countries offer a compelling mix of varying grid structures, renewable energy penetration levels, and electric mobility adoption trajectories, providing a representative testbed for the ODP’s cross-border and multi-sectoral capabilities.
| Parameter | Austria (AT) | Hungary (HU) | Slovenia (SI) | Source/Justification |
| EV stock (2026, thousands) | 343.8 | 100.8 | 50.0 | National Registries, [Statista Research Department, 2023] |
| Annual EV Growth rate (CAGR 2026-35) | NECPs, [BloombergNEF, 2023] | |||
| ET stock (2026, thousands) | 4.4 | 0.4 | 0.3 | Pilot data, [European Automobile Manufacturers Association (2023), ACEA] |
| Annual ET growth rate (CAGR 2026-35) | Industry projections, [McKinsey & Company, 2022] | |||
| RES capacity (Solar + Wind, 2026 GW) | 10.5 | 8.0 | 1.5 | National TSO data, [ENTSO-E, 2022] |
| Annual RES addition (MW/yr avg.) | 664 | 2750 | 150 | NECP targets, [SolarPower Europe, 2023] |
| Avg. electricity retail price (EV, €/kWh) | 0.22 | 0.15 | 0.18 | [Eurostat, 2023], adj. for EV tariffs |
| Grid CO2 intensity (2026, g/kWh) | 120 | 280 | 190 | [European Environment Agency (2023), EEA], projected |
| Value of lost load (VoLL, €/kWh) | 10.0 | 7.5 | 8.0 | Nat. regulator studies, [Agency for the Cooperation of Energy Regulators (2021), ACER] |
| Battery life Ext. (smart charging) | reduction in degradation rate | [Curry et al., 2023, Maheshwari et al., 2020] | ||
| Social cost of carbon (€/tCO2) | Starts at €85 in 2026, rising to €120 by 2035 | EU guidance [European Commission, 2022], linear interpolation. | ||
Table 5 summarizes key country-specific input assumptions for the baseline year 2026 and projected growth rates for the analysis period (details are presented in B). These assumptions are derived from recent national energy and climate plans (NECPs), statistical office data [Eurostat, 2023], reports from international agencies like the IEA [International Energy Agency (2023), IEA], and specialized consultancies (e.g., BloombergNEF, McKinsey) to ensure their credibility and relevance to the respective national contexts. For instance, the EV/ET growth rates reflect ambitious but achievable national targets and industry projections, while RES capacity additions align with NECP goals. Battery-health benefits are quantified using an explicit aging representation within the ODP architecture and monetized as the NPV of the modeled reduction in degradation cost relative to a baseline without ODP. Empirical studies report an 8-12% lower degradation under health-aware charging, narrower state-of-charge windows, C-rate limits, temperature-aware scheduling, and reduced high-SoC dwell, compared with uncoordinated practice [Maheshwari et al., 2020, Curry et al., 2023]. In this study, the empirical range serves as an external validation band lower bound (8%), which is used in the calculation as a conservative figure.
Key benefit categories and their monetization approaches are summarized below. Detailed calculations of each benefit and the specific data required for these calculations are presented in Appendices from A.1.1 to A.1.8.
i) Reduction in operational downtime (ROD) cost: This is from avoided power outages or avoided unserved energy due to improved grid reliability and avoided EV/ET service disruptions. A conservative reduction in baseline outage energy could be avoided due to ODP systems enhancement (1) and AI-driven predictive analytics enables proactive grid management and faster fault restoration. This analysis could be benchmarked against improvements reported in analogous smart grid deployments [U.S. Department of Energy, Office of
Electricity, 2019].
ii) Revenue from optimized energy trading and ancillary services (ROETAS): This is an increase in revenue or cost savings from optimized participation in wholesale energy markets and ancillary services. This benefit is directly influenced by the forecasting accuracy of ODP’s AI models (2).
iii) Cost savings from demand response and peak load reduction (CSDR-PLR): This is the avoided costs from reducing system peak demand through coordinated EV/ET smart charging and DR programs orchestrated by the ODP’s AI. Peak load reductions of (3) are achieved by the ODP’s optimization engine.
iv) Avoided energy curtailment (AEC): It is a monetized value of renewable energy utilized rather than curtailed due to the ODP’s intelligent management. A reduction in baseline curtailment (5) is projected based on simulations where the ODP’s AI (RES forecasting and demand-flexibility matching capabilities) dynamically aligns EV/ET charging with RES availability.
v) Fleet energy and operational savings (FES): This is a reduced energy consumption and operational costs for EV/ET fleets. This includes savings from AI-optimized routing (E-VRPTW algorithms in ODP, yielding an estimated reduction in energy/distance for relevant trips, based on pilot study analogs of AI-optimized fleet routing). A significant component is the monetized value of extended battery lifespan, which is estimated from 8–12% reduction in degradation rate Maheshwari et al. [2020].
This benefit is directly achieved by the ODP’s AI-driven smart charging algorithms, which explicitly incorporate battery health objectives informed by established battery aging models [Maheshwari et al., 2020, Curry et al., 2023].
vi) Grid stability and management cost savings (GSMS): This benefit is considered as reduced costs for DSOs/TSOs associated with grid regulation, voltage support, and congestion management. The benefit is calculated as grid stability savings and grid stability from the forecasting ability of the AI-driven ODP (7). This estimation is based on the ODP’s AI-driven enhanced forecasting and proactive control capabilities, leading to operational efficiency gains analogous to those reported in advanced smart grid operational studies [U.S. Department of Energy, Office of
Electricity, 2019, Electric Power Research Institute
(2020), EPRI] and analyses of AI-driven grid management platforms.
vii) CO2 emission reduction: This is a monetized benefit of greenhouse gas emission reductions attributable to the ODP’s operations, such as increased RES utilization, optimized EV/ET charging due ODP-enabled RES, EV, ET and awareness creation on the use of the operational platform for social cost of carbon reduction.
viii) Reduction in air pollutant (RAP): This is a monetized health and environmental improvement from reduced local air pollutant emissions (10).
The detailed mathematical formulations and parameter derivations for the above benefits are provided in A.1. By summing the discounted annual benefits across the 10-year period, we calculate the present value for each benefit category and the total benefits. Table 6 presents the total present value of benefits over , aggregated across the three countries and categorized by type. Detailed annual benefit projections, underpinning these present values, are also provided in B.
| Benefit category | Component | Present value (€ million) | Share of total benefits (%) |
| Economic benefits | Reduction in operational downtime (ROD) | 70.63 | |
| Optimized energy trading & ancillary services (ROETAS) | 148.20 | ||
| Demand response & peak load reduction (CSDR-PLR) | 170.80 | ||
| Fleet energy & operational savings (FES) | 75.61 | ||
| Avoided energy curtailment (AEC) | 292.99 | ||
| Subtotal economic benefits | 758.23 | 61.5 | |
| Reliability benefitsa | Grid stability & management savings (GSMS)b | 347.99 | |
| Subtotal reliability benefits | 347.99 | 28.2 | |
| Environmental benefits | CO2 emission reduction benefits | 89.07 | |
| Reduction in air pollutant benefits (RAP) | 38.64 | ||
| Subtotal environmental benefits | 127.71 | 10.4 | |
| Total monetized benefits (PV) | 1233.9 | 100.0 |
a
Reliability benefits are substantial. GSMS captures direct grid operational savings. ROD, while categorized under economic, also reflects improved reliability. AEC enhances stability by better integrating variable RES.
b
GSMS: Reduced costs for grid regulation, balancing services (not otherwise captured in ROETAS), and congestion management due to ODP’s predictive control and enhanced visibility.
Economic benefits constitute the largest share ( of total benefits), primarily driven by the AI-enabled optimized energy trading, demand response, and fleet operational savings. Reliability benefits, quantified mainly through grid stability and management savings (GSMS) at ), presents a significant portion of the benefits, indicating the ODP’s potential to enhance grid resilience and potentially defer traditional grid infrastructure investments. Environmental benefits also contribute significantly () highlighting the ODP’s role in accelerating decarbonization.
This enhancement increases the granularity and policy relevance of the environmental assessment without altering the fundamental framework for benefit calculation. Figure 3 presents total benefits by category and their year-by-year evolution. The top-left panel displays the total economic, reliability, and environmental benefits; the top-right panel shows the annual trajectory of each benefit over the project horizon. The bottom panel plots the air-pollutant trends under the existing system and the ODP-enhanced system.
The right panel depicts the yearly trajectory of each benefit in the project horizon. Specifically, country-specific analyses (based on the annual data presented in B and Monte Carlo simulation results in Table 4) reveal nuanced variations in benefit distribution, reflecting national characteristics. For instance, Austria, with its higher initial EV adoption and ambitious electrification targets, sees proportionally greater benefits from fleet energy savings (FES) and a reduction in operational downtime (ROD). Hungary, with its significant RES expansion plans, benefits more substantially from avoided energy curtailment (AEC) and revenue from optimized energy trading and ancillary services (ROETAS) due to the greater flexibility needs. Slovenia, given its smaller scale but gains particularly from optimized ET logistics and enhanced cross-border grid balancing, which the ODP is designed to facilitate. These variations underscore the ODP’s adaptable value proposition across diverse EU member states.
Step 5: Quantify costs
This step involves a comprehensive enumeration and quantification of all costs associated with implementing and operating the ODP over the 2026–2035 period. Costs are categorized into Capital expenditures (CAPEX) and Operational expenditures (OPEX). Detailed breakdowns, differentiating between initial platform development and ongoing integration costs, which are provided in A.2 (CAPEX) and A.3 (OPEX). Below, we explain their main elements:
CAPEX: These are primarily upfront investments incurred during the initial deployment phase (2026) and for system expansion in subsequent years. Major CAPEX components include hardware procurement (sensors, servers, communication devices), core software development for the ODP platform, and crucially, the initial development and training of the AI engine’s specialized models. System integration and project management costs are also included. A significant portion of CAPEX is incremental, related to the integration of new EVs, ETs, and RES installations annually into the ODP-managed ecosystem. Our estimates indicate a total discounted CAPEX of €561.3 million over the 10-year period, with approximately of the initial platform CAPEX directly attributable to AI model development and associated infrastructure (e.g., specialized GPU clusters for training).
OPEX: These are recurring annual costs necessary to operate and maintain the ODP. Key OPEX categories include cloud computing resources for AI inference and data processing, ongoing software maintenance (including continuous AI model monitoring, re-training pipelines to counter data/concept drift, and performance optimization), communication fees, robust cybersecurity operations, labor costs for technical staff (e.g., data scientists, AI engineers, system administrators), and hardware maintenance. Annual OPEX is projected to start at approximately €31.6 million (discounted) in the first year (2026), with an estimated 20% of this directly related to AI operations and maintenance. The total discounted OPEX over 10 years is estimated at €315.9 million.
Figure 4 illustrates the indicative distribution of OPEX in the first year of operation (2026), demonstrating the allocation across various operational aspects, including those supporting AI functionality.
The total discounted investment cost (CAPEX + OPEX) for the ODP deployment and operation over the decade is approximately €877.2 million. Cost estimates are benchmarked against similar smart grid and digital platform deployments in the energy sector [Stromsather et al., 2018, ETIP SNET, 2022] and include reasonable contingencies for unforeseen expenditures. The specific breakdown for AI-related costs captures both initial development (CAPEX) and ongoing operational and maintenance expenses (OPEX), which are essential for sustaining the performance of AI models in a dynamic environment.
CAPEX for shared platform development and core infrastructure, including AI engine development, was allocated to each country based on a composite weighting factor considering projected deployment scale (e.g., number of managed EVs/ETs, RES capacity integrated, and overall energy consumption), as detailed in A.2 and A.3. As a result, Austria accounts for approximately 45% of total costs, Hungary 35%, and Slovenia 20%. Detailed cost breakdowns, including specific required items for AI system development and ongoing AI model maintenance, are comprehensively presented in A.2 and A.3, with annual cost projections provided in B.
Step 6: Compare costs and benefits
In this step, the total present value of monetized benefits is compared against the total present value of costs. Key financial performance indicators, including NPV and BCR, are calculated to assess the overall economic viability of the ODP investment. Table 7 summarizes the aggregated financial outcomes over the 10-year analysis horizon, discounted at 4%.
| Category | Present value (€ million) |
| Total benefits | |
| CAPEX | |
| OPEX | |
| Total costs (CAPEX + OPEX) | |
| NPV (Total benefits Total costs) | 356.7 |
| BCR (Total benefits / Total costs) | 1.41 |
The analysis reveals that total discounted benefits (€1,233.9 million) significantly exceed total discounted costs (€877.2 million). This results in a positive NPV of approximately €356.7 million and a BCR of 1.41. A BCR greater than 1 indicates that the project’s benefits outweigh its costs, with every Euro invested generating €1.41 in benefits. This suggests that the ODP is a financially sound and economically justifiable investment, particularly considering the long-term societal and environmental returns.
Figure 5 provides a visual comparison of the aggregated costs and benefits by their primary categories. Economic benefits constitute the largest share, followed by substantial contributions from reliability and environmental benefits.
The positive NPV and BCR 1 strongly support the economic and financial feasibility of the ODP deployment. These results are based on conservative assumptions for benefit realization and cost estimations, detailed in A.2 and A.3, and include the explicit quantification of AI’s performance-driven contributions.
Step 7: Sensitivity analysis and risk assessment
To assess the robustness of the CBA results to inherent uncertainties in key input parameters, a comprehensive sensitivity and risk analysis was performed. This involved both scenario-based sensitivity testing and probabilistic risk analysis using Monte Carlo simulation. First, scenario-based analyses were conducted by varying critical benefit and cost parameters. Table 8 presents the impact of these variations on the project’s NPV and BCR. Crucially, this includes a specific scenario for AI adoption variability, where the effectiveness factors are reduced.
| Scenario | NPV (€ million) | BCR |
| Base Case | ||
| Total benefits +10% | ||
| Total benefits -10% | ||
| Total costs +10% | ||
| Total costs -10% | ||
| Discount rate 3% (instead of 4% | ||
| Discount rate 5% (instead of 4%) |
The scenario analysis indicates that the ODP investment remains viable (BCR 1, NPV 0) even under reasonably pessimistic assumptions. The sensitivity analysis demonstrates a comfortable margin of safety for the investment. A Monte Carlo simulation ( iterations) was also performed as a probabilistic risk assessment framework. Key parameters such as EV/ET adoption rates, energy price volatility, CAPEX/OPEX unit costs, AI model performance factors (e.g., ranges for forecast accuracy benefits, optimization efficiency gains), and data availability are considered. The specific ranges and distributions were derived from a range of possible real-world outcomes (e.g., slightly lower forecast accuracy than optimal, or less efficient optimization in unforeseen scenarios).
Figure 6 shows the probability distributions of NPV and BCR, providing a probabilistic view of the project’s financial outcomes under uncertainty. The simulations vary energy-price volatility, EV/ET adoption rates, and CAPEX/OPEX around the baseline NPV and BCR estimates.
The Monte Carlo simulation indicates a mean NPV of approximately €365 million and a mean BCR of 1.41. The probability of achieving an NPV greater than zero is estimated to be over , while the probability of a BCR greater than 1.0 is also over . The slight difference between the Monte Carlo mean NPV/BCR and the deterministic base case values can arise from asymmetric probability distributions of input variables, the presence of significant non-linear relationships between input variables and the output metrics, or correlations between input variables, which are inherently captured in a probabilistic approach. These results strongly reinforce the financial robustness of the ODP investment under uncertainty, demonstrating high confidence in its viability even when considering potential variations in adoption and discount rates, as well as other key parameters.
Figure 7 presents the extended sensitivity analysis focusing on four critical parameters: CAPEX, OPEX, data availability, and AI accuracy. The results indicate that NPV ranges from approximately € 100 million to € 700 million, while BCR varies between 1.0 and 1.8. CAPEX assumptions exert the strongest influence on NPV, followed by AI accuracy and data availability, both of which play a crucial role in determining the economic viability of the ODP. These underscore that, although the framework is sensitive to technological performance and data quality, the ODP consistently shows positive value across the explored scenarios.
Qualitative risk assessment identified additional factors: regulatory and policy risks (e.g., changes in market design or data sharing regulations), cybersecurity and data privacy risks (given the sensitive nature of energy and mobility data), technology adoption and integration risks (e.g., slower than expected stakeholder uptake, challenges in integrating with legacy systems), and stakeholder acceptance. Mitigation strategies include proactive engagement with policymakers to shape supportive regulatory frameworks, robust cybersecurity measures that adhere to international standards such as ISO 27001 and GDPR compliance, a phased deployment approach to manage technical complexities, and a user-centric design philosophy to enhance stakeholder acceptance. The sensitivity patterns observed in the aggregated analysis generally hold true at the country level. This suggests that although the absolute values of benefits and costs vary, the project’s overall viability and robustness to uncertainties remain consistent across different national contexts. Monte Carlo simulations performed independently for each country consistently indicate a probability of achieving a BCR 1 exceeding in all three national cases. This further reinforces the economic rationale for the ODP’s deployment at both regional and national scales, providing a high degree of confidence in its financial success despite inherent market and technological uncertainties.
Per-country CBA summary and Monte Carlo simulation (MCS) results
Now, we can see a summary of the CBA outcomes and Monte Carlo simulation (MCS) results for the AI-driven ODP, disaggregated for each of the three countries: Austria (AT), Hungary (HU), and Slovenia (SI). The numerical values presented in Table 9 represent the actual calculated results, based on country-specific assumptions and scaled costs, which demonstrate the national-level impacts derived from the detailed financial and operational models.
| Metric | Austria (AT) | Hungary (HU) | Slovenia (SI) |
| Deterministic results (Base Case) | |||
| Total benefits (PV) | 550.8 | 434.7 | 248.4 |
| Total costs (PV) | 390.4 | 310.8 | 176.0 |
| NPV (PV) | 160.4 | 123.9 | 72.4 |
| BCR | 1.41 | 1.40 | 1.41 |
| Monte Carlo simulation (Mean values from 10,000 Trials) | |||
| Mean NPV (PV) | 168.2 | 127.5 | 74.3 |
| Mean BCR | 1.43 | 1.42 | 1.42 |
| Dominant benefit categories (% of national total benefits) | |||
| Economic (ROD, AEC, FES, ROETAS, CSDR-PLR) | |||
| Environmental (CO2, RAP) | |||
| Reliability (GSMS) | |||
1.
Note: These numerical values are derived from detailed country-specific financial and operational models, reflecting the specific growth rates, market prices, and grid characteristics outlined in Table 5 and other appendices. The slight variations in BCR across countries reflect the unique cost structures and benefit profiles of each national context. Dominant benefit categories highlight the primary drivers of value in each country, influenced by the respective characteristics of their energy and transport systems.
The Monte Carlo simulation results for each country consistently demonstrate a very high probability of positive NPV and BCR greater than 1, reinforcing the robustness of the ODP investment across diverse national contexts within the EU. The distribution of benefits varies slightly, with countries like Austria showing higher absolute economic benefits due to a larger EV/ET stock and higher energy prices, while Hungary benefits significantly from avoided RES curtailment given its ambitious solar expansion. Slovenia, despite its smaller scale, shows strong relative benefits driven by optimized cross-border logistics and effective grid integration.
5 Discussion
5.1 Key findings and insights
The proposed ODP–CBA framework is illustrated with case studies from three countries, but it is general and scalable beyond these settings. While the underlying methodology remains unchanged, differences in sensor unit costs or institutional settings would shift the numerical results but not the validity of the framework itself. The framework’s modular architecture centered on EVs, ETs, RES, and the grid, which remains valid for any jurisdiction, provided basic assumptions and data are available. Scaling to larger systems (e.g., Germany) primarily requires adjusting inputs (fleet sizes, RES capacities, growth projections, grid complexity, sensor unit costs), which change the magnitude of costs and benefits but not the methodology. This design supports both computational scalability and institutional transferability. To this end, the comprehensive CBA conducted for the AI-driven ODP clearly indicates that the proposed platform offers substantial net benefits for the integrated management of electric mobility, renewable energy, and grid operations. The aggregated positive NPV of €356.7 million and a BCR of 1.41 provide a strong economic rationale, signifying that every Euro invested generates approximately €1.41 in benefits over the 10-year analysis period.
The main insight from this study is the substantial economic and operational value unlocked specifically by the ODP’s integrated AI engine. The largest share of the quantified economic benefits stems directly from AI-enabled functionalities, a linkage systematically established through specific ODP performance factors integrated into the benefit monetization (as detailed in Section 4 and A.1). For instance, the revenue from optimized energy trading and ancillary services (ROETAS) and cost savings from demand response and peak load reduction (CSDR-PLR) combined account for over 25.8% of the total benefits, directly attributable to the superior performance of the ODP’s AI forecasting and optimization models.
The case studies across Austria, Hungary, and Slovenia demonstrate the ODP’s adaptability and value proposition across varied national energy and transport landscapes. The substantial reliability benefits (€348 million in GSMS, alongside €171 million in CSDR-PLR as an economic benefit) underscore the ODP’s potential as a highly effective non-wires alternative, achieved through improved system visibility, AI-driven proactive control, and enhanced coordination across sectors. From a societal and policy perspective, the significant environmental benefits (€127 million from CO2 and air pollutant reductions) are critical, highlighting the ODP’s role in accelerating decarbonization. The robustness of AI models to real-world conditions, such as data drift, concept drift, and unexpected events, is planned to be managed through continuous performance monitoring, automated retraining pipelines, and robust data validation, all of which are factored into the recurring OPEX (A.3), ensuring sustained long-term benefits.
In addition to the aggregated results, Table 9 presents the country-specific CBA outcomes and Monte Carlo simulations for Austria, Hungary, and Slovenia. This disaggregation highlights national variations in benefits and costs due to grid structure, RES penetration, and policy context. The results confirm that economic benefits are the dominant category in all three countries, while reliability and environmental benefits contribute proportionally more in Hungary and Slovenia. Across all countries, the benefit–cost ratio remains above one under the 90% confidence interval, demonstrating the robustness of the investment case even under country-specific risk and uncertainty.
5.2 Limitations and implementation challenges
While the comprehensive CBA provides a strong economic rationale, several limitations warrant acknowledgment. First, while rigorous validation were performed, and simulations incorporate realistic assumptions, there remains a potential for variance in real-world AI model performance versus simulated benchmarks, particularly under unforeseen extreme conditions. Future pilot projects are essential for further empirical validation. Second, the optimal functioning of the ODP’s AI engine relies on sustained, high-quality, and comprehensive data availability from a multitude of disparate sources. These include real-time EV and ET fleet telemetry, smart charging infrastructure, renewable energy forecasts, market data from EPEX Spot and ENTSO-E, weather models, and public datasets from Eurostat and the EEA. Third, our CBA does not include primary behavioral studies (e.g., discrete-choice experiments or user trials), beyond stakeholder engagement through interviews and workshops. Although behavioral factors are incorporated indirectly via the stakeholder-informed multi-criteria assessment and sensitivity analyses on adoption and usage, targeted behavioral studies to estimate adoption elasticities, routing/charging preferences, and data-sharing incentives would further strengthen the evidentiary basis and validity of the CBA. Four, Although a 10-year window is appropriate for a digital platform with sub-decadal upgrade cycles, some grid-adjacent hardware may realize benefits over longer periods. Evaluating selected components over 20–25 years—and reporting horizon sensitivities—can further validate long-run value in infrastructure-heavy contexts.
In addition, the following three prevalent precautions should be considered further considered:
Valuation and behavioral uncertainty: Some societal benefits (e.g., improved air quality, user convenience, consumer empowerment) are not fully monetized, which may understate ODP’s total value. In addition, the analysis assumes rational uptake of incentives by EV/ET users and fleets; however, real-world participation and rebound effects may differ, necessitating the use of behavioral modeling and longitudinal evidence.
Technical robustness and data risk: Large-scale data exchange and cloud operation introduce cybersecurity and governance risks. Model accuracy can also drift as mobility patterns, RES penetration, and technology costs evolve. Mitigation requires strong data governance frameworks (including privacy-preserving analytics and independent audits), continuous retraining and performance monitoring, and operational fallbacks. All these evaluations could be better estimated if there were already operating ODPs, even on a small scale.
Regulatory and institutional fit: Scalability depends on alignment with grid codes, tariff design, and data-sharing rules. Fragmented or evolving policies can increase coordination costs and slow deployment; staged pilots and regulatory sandboxes can help mitigate implementation risks and harmonize standards. That said, qualitative stakeholder analyses may not fully capture behavioral and adoption challenges. Empirically estimating the costs and benefits of these elements during implementation could enhance the reliability of the CBA framework.
6 Conclusion
This paper presents a comprehensive cost-benefit analysis of an AI-driven operational digital platform developed for the integrated management of electric vehicles, electric trucks, renewable energy sources, and the power grid. The CBA framework was designed in alignment with European Commission guidelines for energy infrastructure projects and tailored to explicitly capture AI-specific value streams. These include improvements in forecasting accuracy, operational optimization, grid reliability, and emissions reduction, each of which is monetized using transparent and reproducible methodologies. Importantly, the framework incorporates both capital and operational expenditures, particularly those linked to AI model development, deployment, and maintenance, alongside benefits categorized as economic, reliability, and environmental.
The analysis shows that the proposed AI-driven ODP delivers substantial economic and financial value. Over the 2026–2035 operating period, it achieves a projected net present value of about €356.7 million and a benefit–cost ratio of 1.41, indicating a strong investment case. These outcomes are driven by the platform’s multi-layered architecture; its AI core—combining optimization and machine-learning algorithms—is a critical value enabler. The ODP improves energy and operational efficiency, enables predictive congestion mitigation, optimizes demand response, and reduces battery degradation through health-aware charging strategies. Using a dedicated framework, the CBA systematically monetized both benefits and investment/operating costs. Robustness was confirmed through sensitivity analyses and Monte Carlo simulations.
To validate cross-border and cross-sector applicability, the CBA was applied to Austria, Hungary, and Slovenia—each with distinct energy and transport profiles. The results illustrate how national characteristics shape the distribution of benefits. At the country level, Austria achieves substantial fleet-efficiency gains and extended battery-life benefits; Hungary sees improved RES integration and reduced curtailment; and Slovenia has a comparative advantage in grid flexibility and optimized ET logistics. These variations underscore the ODP’s scalability and adaptability across diverse regulatory, technical, and market contexts.
Beyond financial metrics, the study offers strategic insights. For policymakers, the results confirm the crucial role of advanced digital infrastructure in achieving EU-wide decarbonization goals under REPowerEU and the Fit for 55 initiatives. The ODP presents a viable pathway to accelerate transport electrification, enhance grid reliability, and reduce emissions through AI-based coordination. For industry stakeholders, including DSOs, TSOs, fleet operators, and aggregators, the analysis provides a quantitative basis for investing in interoperable, intelligent systems capable of delivering both operational gains and new service opportunities.
At the same time, the study acknowledges key limitations. The optimal functioning of the AI engine depends on sustained access to high-quality, heterogeneous data sources, such as real-time telemetry from EVs and ETs, smart charging infrastructure, market platforms (e.g., EPEX Spot, ENTSO-E), and weather models. Moreover, generalizing the results to other countries requires careful consideration of differences in data infrastructure, policy environments, and electrification maturity. Additionally, intangible benefits, such as improved public health from air quality improvements or increased user convenience, were not monetized, likely resulting in a conservative estimate of the platform’s true societal value.
In conclusion, the integrated management of EVs, ETs, RES, and the grid via AI-driven digital platforms is shown to be not only technically feasible but also economically and environmentally beneficial. The proposed ODP offers a compelling blueprint for next-generation energy–mobility convergence, supporting the emergence of smart, flexible, and low-carbon energy ecosystems. By holistically orchestrating distributed assets through AI models, the platform bridges the gap between high-level policy goals and on-the-ground operational needs.
Future work should focus on several key areas to further validate and expand the ODP’s capabilities and impact. This includes the initiation of pilot projects for empirical validation of the ODP’s AI models and overall system performance in real-world operational environments, providing concrete data on benefits and operational challenges. A deeper exploration of advanced V2X (Vehicle-to-Everything) services, including vehicle-to-home (V2H) and vehicle-to-grid (V2G) applications, could unlock further synergies beyond ancillary services. Detailed distribution grid impact modeling, specifically for local voltage stability and transformer loading under high EV penetration and dynamic charging, is crucial. Research into advanced cybersecurity measures tailored for AI-driven energy platforms, including AI-specific threat detection and mitigation, is paramount given the critical infrastructure nature. Policy and market design innovation are essential to incentivize broad stakeholder participation and ensure fair value distribution across the ecosystem. Finally, extending the ODP concept to integrate other flexible resources such as smart heat pumps, stationary battery storage, and industrial flexible loads could further enhance system-wide optimization and accelerate decarbonization efforts.
Acknowledgements
This research was supported by the BEGONIA project, which has received funding from the European Commission under grant agreement No. 01133306.
Appendix A Detailed cost-benefit analysis methodology and data
This appendix provides further details on the methodologies and data underpinning the CBA presented in this paper, ensuring transparency and reproducibility of the quantitative analysis.
A.1 Benefit monetization formulas and key parameters
This section presents the primary formulas and key parameters used for monetizing benefits using CBA framework. All monetary values are expressed in real terms (Euro 2025), and future cash flows are discounted to their present value using an economic discount rate of 4%. Each benefit’s realization is linked to the capabilities of the AI engine described in the ODP architecture. In the following, we present the formulas for calculating each benefit.
A.1.1 Reduction in operational downtime (ROD) cost
The benefit from reduced operational downtime is calculated as the monetized value of avoided unserved energy due to improved grid reliability and faster fault recovery facilitated by the ODP’s AI-driven predictive maintenance and enhanced situational awareness. The benefit from reduced operational downtime is modeled as:
where is the baseline operational downtime cost, is the energy self‐efficiency improvement fraction, and is the ODP‐enabled downtime cost savings, reflecting faster fault recovery and improved grid reliability due to AI‐driven predictive maintenance and situational awareness. The product captures the adjusted downtime cost under improved self‐efficiency, and multiplying by converts this into the annual benefit from avoided unserved energy, consistent with analogous smart‐grid reliability improvements [U.S. Department of Energy, Office of Electricity, 2019]. Parameter values are presented in Table 10.
A.1.2 Revenue from optimized energy trading and ancillary services (ROETAS)
This benefit captures the incremental revenues and cost savings enabled by the ODP’s AI forecasting and optimization across energy arbitrage and ancillary-service market:
where:
1.
and are the hourly energy–arbitrage revenues (in €) in hour with and without ODP, respectively. We compute hourly arbitrage revenue as
with the flexible energy shifted in hour (MWh), the realized buy/sell price spread (€/MWh), and the round-trip efficiency.
2.
denotes the set of ancillary-service products. For each and hour , and are the ODP’s hourly revenue and cost in €; and are the corresponding baseline values.
If one prefers a compact proxy for arbitrage gains due to efficiency/forecasting improvements, the first term in 2 can be approximated by
where is the net uplift fraction attributable to ODP (dimensionless), is the assumed average electricity price (€/MWh), is baseline annual renewable output (MWh) with installed capacity (MW) and full-load hours (h/yr), and is baseline annual EV+ET energy (MWh) given aggregate charging capacity (MW) and h/yr.
A.1.3 Cost savings from demand response and peak load reduction (CSDR‐PLR)
Avoided costs from reducing system peak demand are realized through ODP’s AI‐driven demand response (DR) and peak‐load‐shifting capabilities. We model the annual benefit as
where the first term represents the benefits from cost savingsresulting from demand response, and the second term represents the benefits from peakload revenue. [€/MWh] is the unit cost saving from demand response (CSDR) per shifted MWh; [MW] is the installed RES capacity and [h/yr] its full‐load hours, hence is the annual RES‐generation baseline in MWh. The factor is the fraction of RES output that can be shifted via ODP-enabled. Similarly, [MW] is the aggregated EV charging capacity, [MW] is the total ET charging capacity, the terms and are the annual EV and ET baseline energy (MWh) multiplied by the shiftable fractions and due to ODP.
In the second term, [€/MWh] denotes the peak‐load reduction (PLR) revenue per shifted MWh, and is the fraction of total annual energy that can be shifted away from peak hours. We also have:
where is the annual volume (MWh) shifted out of peak hours. Thus, the second term represents revenue from peak‐load reduction at peak‐period prices.
A.1.4 Avoided energy curtailment (AEC)
Monetized value of RES utilized rather than curtailed, enabled by ODP’s AI‐driven RES forecasting and flexible load matching. This benefit is calculated using the following formula:
where is the fraction of RES curtailment avoided, [MW] is the installed RES capacity, and [€/MWh] is the unit cost associated with avoided curtailment. Thus, gives the annual baseline RES generation (MWh), multiplying by yields the avoided MWh, and multiplying by converts to euros.
A.1.5 Fleet energy and operational savings (FES)
Reduced EV/ET operational costs are achieved through ODP’s AI‐optimized routing and charging, which extend battery lifespan and improve energy efficiency. The total annual benefit is calculated as follows:
where is the overall ODP‐enabled savings fraction, and are the total number of EVs and ETs, respectively, and are the effective efficiency factors capturing reduced energy consumption and extended battery life, and and are the annual per‐vehicle energy and battery‐degradation savings under ODP. The factor consolidates multiple AI‐driven improvements: battery‐health optimization (reducing degradation by 8–12% via SoC and C‐rate management [Maheshwari et al., 2020, Curry et al., 2023]) and routing efficiency gains (3–5% savings from E-VRPTW optimization). Thus, reflects a conservative 1% net reduction in total fleet operating cost when multiplying the baseline EV and ET savings terms.
A.1.6 Grid stability and management cost savings (GSMS)
Reduced DSO/TSO costs for grid stability arise from several ODP‐enabled improvements (Section 3.1). Grid congestion and reliability benefits are quantified through reductions in expected energy not served and congestion events. Baseline reliability is estimated from national outage statistics and grid development reports for Austria, Hungary, and Slovenia Statista Research Department [2023]. The ODP’s impact is modeled as the incremental reduction in expected energy not served, reflecting improved flexibility and RES integration. Monetization is based on country-specific Values of Lost Load, as reported in regulatory and market studies. Thus, we calculate the annual benefit of GSMS as
where is the ODP‐enabled forecasting improvement fraction (informed by smart‐grid pilot results [U.S. Department of Energy, Office of Electricity, 2019, Electric Power Research Institute (2020), EPRI]), is the total number of EVs, is the total number of ETs, [MWh/yr] and [MWh/year] are the annual charging energy per EV and per ET, respectively, and [€/MWh] is the unit fuel‐saving value. The product captures savings from reduced fuel use due to improved timing of EV/ET charging. The term represents avoided operational‐downtime costs: is the baseline downtime cost, [MW] is installed RES capacity, and is the RES full‐load hours, so that multiplying by estimates the portion of downtime cost deferral attributable to RES‐driven stability improvements.
Finally, is the ODP‐enabled grid‐stability savings fraction. Multiplying by where [€/MWh] is the market‐price reduction rate, yields additional benefits from reduced imbalance and congestion charges when EV/ET charging shifts under ODP’s real‐time control.
A.1.7 CO2 emission reduction benefits
Monetized GHG emission reductions arise from ODP’s AI‐driven increase in RES utilization and optimized EV/ET charging. To increase granularity, we now differentiate both temporally and spatially. We model the annual benefit as
where is the ODP‐enabled reduction fraction, [European Commission, 2022]. where is the ODP‐enabled reduction fraction varying by time and country, is the number of EVs in operation in country at time and [MWh/vehicle] is their average energy demand after AI‐optimized charging. Similarly, is the number of ETs, and [MWh/vehicle] is their annualized energy demand. denotes the renewable generation (capacity multiplied by capacity factor) available in country at time . The factor [tonne CO2/MWh] is the hourly marginal emission factor, and [€/tonne CO2] is the country‐specific social cost of carbon. Thus, the summation over and captures the total avoided CO2 emissions, weighted by the temporally varying marginal emission intensity and regionally differentiated damage valuation.
A.1.8 Reduction in air pollutant (RAP)
Monetized health and environmental benefits from reduced local air pollutants arise when ODP’s AI enables cleaner generation, demand response, and transport.
Monetized health and environmental benefits from reduced local air pollutants arise when ODP’s AI enables cleaner generation, demand response, and optimized transport. We model this benefit as
where [kg/km] and [kg/km] are pollutant emission factors for EVs and ETs in country at time , reflecting local fleet composition and technology. and [km/year] are annual driving distances per EV and ET, and and are the respective vehicle counts. Thus, and give the pollutant reductions (kg) attributable to AI-enabled charging and scheduling.
The term [kg/MWh] denotes the hourly marginal emission factor for pollutant , while , , captures the RES output shifted and decarbonized through ODP optimization.
Finally, [€/kg] represents the country-specific marginal damage cost of pollutant , accounting for local air quality, health impacts, and income levels. The summation over and ensures that spatial and temporal heterogeneity in exposure and generation mix is reflected in the monetized benefit assessment, improving granularity and policy relevance relative to static average values.
| Name | Symbol | Unit | Value | Source |
| Baseline downtime cost | M€/yr | National statistics. | ||
| Self-efficiency improvement | % | Assumption (tested in sensitivity). | ||
| ODP downtime savings rate | % | Assumption. | ||
| Avg. electricity price | €/MWh | ENTSO-E [2025]. | ||
| Arbitrage energy share | % of energy | Assumption (proxy uplift). | ||
| RES capacity | MW | ENTSO-E [2025]. | ||
| RES full-load hours | h/yr | Eurostat [2025], European Environment Agency [2024]. | ||
| EV charging capacity | MW | dataset (Table 14); national stock data per Eurostat Eurostat [2025]. | ||
| ET charging capacity | MW | dataset (Table 14). | ||
| DR unit saving | €/MWh | National regulator tariff schedules. | ||
| Peak-shiftable share | % of energy | Assumption. | ||
| Avoided curtailment share | % of RES | Assumption. | ||
| Curtailment valuation | €/MWh | Assumption (opportunity cost proxy). | ||
| Fleet savings factor | % | Conservative assumption. | ||
| Per-EV annual saving | €/vehyr | IEA Global EV Outlook International Energy Agency [2024]. | ||
| Per-ET annual saving | €/vehyr | Rodríguez et al. [2023]. | ||
| Grid-stability improvement | % | Badanjak and Pandžić [2021]. | ||
| Fuel value (avoided) | €/MWh | ACER and CEER [2024]. | ||
| Grid-stability savings rate | % | Assumption. | ||
| Market price (stability) | €/MWh | ACER and CEER [2024]. | ||
| Value of lost load | €/MWh | CEPA study for ACER Cambridge Economic Policy Associates (2018) [CEPA]. | ||
| CO2 reduction factor | % | Assumption. | ||
| Emission factor (grid) | tCO2/MWh | European Environment Agency [2024]. | ||
| CO2 price | €/tCO2 | European Energy Exchange (2025) [EEX]. | ||
| Pollutant emission factor | kg/MWh | EEA and EMEP/CLRTAP [2019]. | ||
| Pollutant unit cost | €/kg | European Environment Agency [2020]. |
A.2 CAPEX breakdown
This appendix details the capital expenditure (CAPEX) components for the AI-driven ODP. The figures in Table 11 represent the discounted present value (PV) of costs over a ten-year operational period (2026-2035), aggregated for deployment across Austria, Hungary, and Slovenia. The breakdown distinguishes between ongoing hardware investments for expanding sensor networks (which are incremental annually) and initial one-time investments for the core technology platform. The AI-specific CAPEX includes initial development of AI models and the dedicated computing infrastructure required for their training and deployment.
| Cost category | Total (PV) |
| Hardware (sensors and edge devices - incremental) | |
| Incremental sensors for new EVs | 280.00 |
| Incremental sensors for new ETs | 120.00 |
| Incremental sensors for new RES installations | 71.65 |
| Subtotal hardware (incremental) | 471.65 |
| Core technology platform (initial investments and upgrades) | |
| Adaptive environmental sensor arrays (initial deployment) | 20.00 |
| Weather forecasting sensor infrastructure | 1.20 |
| Connectivity infrastructure (5G/LPWAN) | 3.00 |
| Communication protocols | 3.00 |
| Message queue systems (Kafka) | 3.00 |
| Databases and data lakes (setup) | 3.00 |
| Data integration middleware (FIWARE Context Broker) | 4.50 |
| API management gateway | 4.50 |
| Market integration middleware (EPEX SPOT APIs) | 4.50 |
| AI models for grid optimization (development and initial training) | 15.00 |
| Demand response tools (integration) | 6.00 |
| Route optimization engines (development and integration) | 9.00 |
| Weather forecasting algorithms (development and integration) | 3.00 |
| End-user mobile application development | 0.90 |
| Charging navigation systems integration | 3.00 |
| Comprehensive analytics tools development | 3.00 |
| Administrative dashboard development | 3.00 |
| Software deployment and system integration | 3.00 |
| Subtotal core technology platform | 89.60 |
| Total CAPEX (PV) | 561.25 |
1.
Note: Incremental sensor CAPEX reflects the cumulative discounted cost of equipping newly adopted EVs, ETs, and RES units annually with ODP-compatible sensors and edge devices. Core technology platform costs represent foundational one-time investments for the ODP core infrastructure and software components, including dedicated AI development. These figures are based on industry benchmarks for similar digital platform deployments in the energy sector, scaled for the three-country scope. Contingencies are embedded within these estimates.
A.3 OPEX breakdown
The figures in Table 12 represent the discounted present value (PV) of recurring annual costs OPEX over a ten-year operational period (2026-2035), aggregated for deployment across Austria, Hungary, and Slovenia. The Annual average PV provides an annualized perspective of these costs, calculated from the total 10-year discounted OPEX. Costs specifically related to sustaining and evolving the AI engine are explicitly considered.
| Cost category | annual Avg. PV |
| Cloud computing (infrastructure, VMs, storage, networking) | 3.9 |
| AI inference and data processing (real-time predictions, model serving) | 3.9 |
| Software maintenance and updates (platform, third-party libraries) | 2.2 |
| AI model maintenance, retraining, and performance monitoring | 2.5 |
| Grid and sensor maintenance (upkeep of physical assets) | 2.8 |
| Cybersecurity and threat monitoring (platform-wide) | 3.6 |
| Customer support and data analytics (stakeholders and users) | 1.9 |
| TSO/DSO and energy market integration (data feeds, API costs) | 2.8 |
| Grid expansion and new EV/ET integrations (operational, excl. CAPEX) | 3.2 |
| Network maintenance (5G/LPWAN) | 3.0 |
| Market trading fees (transactions, commissions) | 2.4 |
| Miscellaneous costs (training, compliance, licenses) | 2.8 |
| Labor costs (technical staff, data scientists, operational support) | 3.0 |
| Total annual average OPEX (PV) | 31.6 |
| Total 10-Year OPEX (PV) | 315.9 |
1.
Note: The Annual Average PV is derived from the total 10-year discounted OPEX. The distribution across categories is proportionally guided by detailed first-year operational cost assessments, encompassing all recurring expenses for platform operation, maintenance, and support. AI-specific OPEX (e.g., Cloud computing for AI inference, AI model maintenance) accounts for approximately 20% of total OPEX, crucial for sustaining the AI engine’s performance. Contingencies are embedded within these estimates.
To increase transparency in how AI-related expenditures are represented, we distinguish between one-time CAPEX (Table 11) and recurring OPEX (Table 12). CAPEX categories capture the upfront development and integration of AI modules, while OPEX categories cover recurring costs such as cloud hosting, AI inference and retraining, cybersecurity, and customer analytics. This breakdown clarifies that AI-specific costs are explicitly embedded in both cost categories, consistent with procurement benchmarks from AWS, Azure, Eurostat ICT reports, and IEA/ACER market integration studies. Table 13 summarizes the unit assumptions for AI model training, cloud infrastructure, and maintenance. These values are based on European and international benchmarks and were scaled using adoption and growth assumptions presented in Appendix A2. Together, they account for approximately 20% of total OPEX and 15% of total CAPEX.
| AI function / resource | Unit cost assumption | Source |
| AI model training (grid + transport) | €1-1.2 million per comprehensive model | [AWS, 2024, Azu, 2024] |
| AI inference (real-time predictions) | €0.05–0.07 per CPU-hour | [AWS, 2024, Azu, 2024] |
| Cloud storage and integration | €0.02–0.03 per GB-month | [Eurostat, 2023] |
| Model retraining and monitoring | €1.2–1.5M per model-year | [European Commission, 2023, International Energy Agency, 2023] |
| Cybersecurity and threat monitoring | €200–300k per country-year | [International Energy Agency, 2023] |
| API/market integration | €1.0–1.3M per country-year | [Agency for the Cooperation of Energy Regulators (2023), ACER] |
| Labor (AI/data engineers, ops staff) | €45–60k per FTE per year | [Eurostat, 2023] |
Appendix B Annual projections and underlying data for the ODP case studies
This appendix provides a detailed overview of the annual projections for key parameters, monetized benefits, and costs specifically for the AI-driven ODP case studies across Austria, Hungary, and Slovenia. These projections form the basis for the discounted present values presented in the main paper and are derived from the comprehensive modeling framework and assumptions outlined in Section 3.
B.1 Annual EV/ET Stock and RES capacity projections
The adoption rates of EVs and ETs, along with the growth of RES capacity, are fundamental drivers of the ODP’s benefits and costs. Table 14 and Table 15 present the annual projected stock of EVs and ETs, and the installed RES capacity for each of the three countries over the 10-year analysis period (2026-2035). These projections are consistent with national energy and climate plans (NECPs), established industry forecasts, and historical trends, as referenced in Table 5.
| Year | EV Stock (thousands) | ET stock (thousands) | ||||
| Austria | Hungary | Slovenia | Austria | Hungary | Slovenia | |
| 2026 | 343.8 | 100.8 | 50.0 | 4.4 | 0.4 | 0.3 |
| 2027 | 398.8 | 118.8 | 57.7 | 4.9 | 0.5 | 0.3 |
| 2028 | 462.6 | 140.0 | 66.6 | 5.5 | 0.5 | 0.4 |
| 2029 | 536.6 | 165.1 | 76.9 | 6.2 | 0.6 | 0.5 |
| 2030 | 622.5 | 194.7 | 88.8 | 7.0 | 0.7 | 0.5 |
| 2031 | 722.1 | 229.6 | 102.6 | 7.9 | 0.8 | 0.6 |
| 2032 | 837.7 | 270.6 | 118.5 | 8.9 | 0.9 | 0.7 |
| 2033 | 971.7 | 318.9 | 136.9 | 10.1 | 1.0 | 0.8 |
| 2034 | 1127.2 | 375.9 | 158.2 | 11.4 | 1.1 | 0.9 |
| 2035 | 1307.5 | 443.3 | 182.7 | 12.8 | 1.3 | 1.0 |
| Year | Austria (GW) | Hungary (GW) | Slovenia (GW) |
| 2026 | 10.5 | 8.0 | 1.5 |
| 2027 | 11.2 | 10.7 | 1.6 |
| 2028 | 11.8 | 13.5 | 1.8 |
| 2029 | 12.5 | 16.3 | 1.9 |
| 2030 | 13.2 | 19.1 | 2.1 |
| 2031 | 13.9 | 21.9 | 2.2 |
| 2032 | 14.6 | 24.7 | 2.4 |
| 2033 | 15.3 | 27.5 | 2.5 |
| 2034 | 16.0 | 30.3 | 2.7 |
| 2035 | 16.7 | 33.1 | 2.8 |
B.2 Annual monetized benefits of the ODP (aggregated)
Table 16 presents the annual monetized benefits of the ODP, aggregated across Austria, Hungary, and Slovenia, prior to discounting. These figures demonstrate the year-on-year growth of benefits as the ODP’s functionalities mature and as EV/ET adoption and RES integration increase.
| Year | ROD | ROETAS | CSDR-PLR | FES | AEC | GSMS | CO2 | RAP | Total annual benefits |
| 2026 | |||||||||
| 2027 | |||||||||
| 2028 | |||||||||
| 2029 | |||||||||
| 2030 | |||||||||
| 2031 | |||||||||
| 2032 | |||||||||
| 2033 | |||||||||
| 2034 | |||||||||
| 2035 | |||||||||
| Total |
B.3 Annual costs of the ODP (aggregated)
Table 17 details the annual CAPEX and OPEX for the ODP, aggregated across Austria, Hungary, and Slovenia, prior to discounting. These figures illustrate the investment profile over the 10-year period, with initial higher CAPEX followed by recurring operational costs.
| Year | Annual CAPEX (€ million) | Annual OPEX (€ million) | Total annual costs (€ million) |
| 2026 | |||
| 2027 | |||
| 2028 | |||
| 2029 | |||
| 2030 | |||
| 2031 | |||
| 2032 | |||
| 2033 | |||
| 2034 | |||
| 2035 | |||
| Total |
1.
Note: Annual CAPEX and OPEX are discounted sums over the 2026–2035 period, aggregated across Austria, Hungary, and Slovenia. The Total row shows the ten‐year aggregate costs for each category. In addition, a one‐time upfront fixed cost of €89.6 million is incurred in 2026 to cover connectivity (5G/LPWAN), communication protocols and message queues, databases and data lakes, data integration middleware, market integration middleware, AI models for grid optimization, demand response tools, route optimization engines, weather-forecasting algorithms, software deployment and integration.